### v1

# Morteza Rahimi    Susan Ghaderi    Yves Moreau    Masoud Ahookhosh

Abstract
This paper is devoted to the class of paraconvex functions and presents some of its fundamental properties, characterization, and examples that can be used for their recognition and optimization. Next, the convergence analysis of the projected subgradient methods with several step-sizes (i.e., constant, nonsummable, square-summable but not summable, geometrically decaying, and Scaled Polyak’s step-sizes) to global minima for this class of functions is studied. In particular, the convergence rate of the proposed methods is investigated under paraconvexity and the Hölderian error bound condition, where the latter is an extension of the classical error bound condition. The preliminary numerical experiments on several robust low-rank matrix recovery problems (i.e., robust matrix completion, image inpainting, robust nonnegative matrix factorization, robust matrix compression) indicate promising behavior for these projected subgradient methods, validating our theoretical foundations.
Keywords:
Nonsmooth nonconvex optimization, Paraconvex functions, Projected subgradient methods, Hölderian error bound, Linear convergence, Robust low-rank matrix recovery.
MSC:
90C06, 90C25, 90C26, 49J52, 49J53.
1 Introduction
Subgradient methods and their projected variants were originally introduced for solving convex nonsmooth problems, dating back to 1960, see, for example, ermol1966methods ; polyak1967general ; polyak1987introduction ; shor2012minimization . Due to the simple structure and low memory requirement, subgradient methods have been employed increasingly, nesterov2009primal ; nesterov2014subgradient ; nesterov2024primal , especially when the objective function does not have a simple structure for splitting algorithms; cf. ahookhosh2018optimal ; ahookhosh2017optimal ; ahookhosh2018solving ; neumaier2016osga and references therein. Although there are many interesting convex applications, the recent trends in application domains, such as signal and image processing, machine learning, data science, statistics, sparse optimization, and so on, are introducing many nonsmooth and nonconvex functions without simple structures for splitting algorithms. These applications involve, but are not restricted to, image deblurring liu2021efficient , image inpainting yu2019new , compressed sensing chartrand2008restricted , sparse recovery wen2018survey , matrix factorization chi2019nonconvex , nonnegative matrix factorization gillis2020nonnegative , matrix completion sun2016guaranteed , tensor factorization welling2001positive , phase retrieval eldar2014phase , deep neural networks jain2017non , and so on. This motivates the quest for developing efficient subgradient methods for such large to huge-scale problems, which is the main objective of the current work.
In the current paper, we are interested in solving the nonsmooth and nonconvex optimization problem
$\min_{x\in X}f(x),$ (1.1)
where the function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}=\mathbb{R}\cup\{+\infty\}$ is proper, nonsmooth, and nonconvex, and $X\subseteq\mathbb{R}^{n}$ is a nonempty, closed, and convex set representing our feasible set. We assume that $\mathcal{X}^{*}$, the set of optimal solutions, is nonempty and $f^{*}>-\infty$ is the optimal value. As opposed to the convex cases, solving nonsmooth and nonconvex problems can be much more complex, considering the following reasonings: (i) the lack of differentiability for the objective function requires the nonemptiness of (Fréchet/Mordukhovich/Clarke) subdifferentials of $f$ denoted as $\partial f(\cdot)$, cf. clarke1990optimization ; (ii) the landscape of nonconvex function can be considerably more complex involving minima, maxima, and saddle points; (iii) the local minimizers are no longer the global ones. On top of these difficulties, the number of variables for application problems of the form (1.1) can be huge, i.e., any algorithm for dealing with such applications should have a simple structure.
The first-order methods are among the suitable schemes to handle the optimization problem (1.1), relying only on the objective values, the gradient (in the smooth case) or subgradient (in the nonsmooth case) of the underlying function bertsekas2003convex ; nesterov2018lectures ; ruder2016overview . In order to handle large- to huge-scale problems, it is better to avoid the computation of objective functions. To this end, there are two main strategies: (i) the constant step-size for gradient-based methods requiring $f\in\mathcal{C}^{1,1}_{L}$ nesterov2018lectures or $f$ to be relatively smooth bauschke2016descent ; lu2018relatively ; ghaderi2024nineuclidean ; (ii) the predefined step-sizes for descent first-order methods nesterov2024primal ; davis2018subgradient . Among all other first-order methods, we are interested in projected subgradient method given by
${\mathopen{}\left\{\begin{array}[]{l}\text{Select}\leavevmode\nobreak\ \,\zeta_{k}\in\partial f(x_{k}),\vspace{2mm}\\ \text{Set}\,x_{k+1}=\operatorname{proj}_{X}{\mathopen{}\left(x_{k}+\alpha_{k}\tfrac{\zeta_{k}}{\|\zeta_{k}\|}\right)\mathclose{}},\end{array}\right.\mathclose{}}$
where $x_{k}$ is the current iteration, $\alpha_{k}>0$ is step-size, and the operator $\operatorname{proj}_{X}$ represents the projection onto $X$. This algorithm does not require any function values and has the step-size predefined before its implementation, making it suitable for large-to huge-scale problems. The subgradient-based methods have received much attention in convex setting auslender2009projected ; bertsekas2003convex ; johnstone2020faster ; kiwiel2004convergence ; nedic2001incremental ; nesterov2009primal ; nesterov2018dual ; shor2012minimization . In the nonconvex setting, most studies of subgradient methods have only been carried out for limited classes of functions such as quasiconvex functions kiwiel2001convergence ; hu2020convergence ; hu2015inexact ; quiroz2015inexact and weakly convex functions davis2018subgradient ; li2021weakly , along with references therein. This is our primary motivation to explore the convergence analysis of subgradient methods for larger classes of functions.
We here focus on the class of $\nu$-paraconvex functions $\nu\in(0,1]$. A function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is named $\nu$-paraconvex if there exists some $\rho\geq 0$ such that
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+\rho\min\{\lambda,1-\lambda\}\|x-y\|^{1+\nu},$ (1.2)
for each $x,y\in X$ and $\lambda\in[0,1]$. The class of $\nu$-paraconvex functions were first introduced by Rolewicz rolewicz1979paraconvex ; rolewicz1979gamaparaconvex for a class of multi-valued mappings. In particular, Rolewicz rolewicz1979paraconvex demonstrated that $1$-paraconvex functions can be expressed as the difference between a convex function and a quadratic term, where these functions are also known as weakly convex functions vial1983strong satisfying the property that the perturbed function $x\mapsto f(x)+\rho\|x\|^{2}$ is convex for some $\rho\geq 0$. However, Rolewicz rolewicz1979gamaparaconvex provided a counterexample showing that this property does not hold for $\nu\in(0,1)$. Nevertheless, we establish that if the function $x\mapsto f(x)+\rho\|x\|^{1+\nu}$ is convex for some $\rho\geq 0$, then $f$ qualifies as $\nu$-paraconvex. Thus, the class of $\nu$-paraconvex functions extends beyond the class of weakly convex functions, providing a broader framework for analyzing nonconvex problems. Additional characterizations and properties (like continuity, local Lipschitzness, subdifferentiability, subdifferential monotonicity, and so on) of these functions were discussed in jourani1996open ; jourani1996subdifferentiability ; van2008paraconvex . Moreover, these functions arise in many applications such as optimal control, image processing, machine learning, compressed sensing, and so on; cf. cannarsa2004semiconcave ; li2021weakly ; davis2018subgradient .
In absence of the strong convexity of the objective function, studying the convergence rate of optimization methods under variants of error bound conditions has a long story in optimization; see, e.g., bolte2007lojasiewicz ; bolte2017error ; burke2002weak ; burke1993weak ; davis2018subgradient ; ferris1991finite ; johnstone2020faster ; li2013global ; luo1993error ; pang1997error ; xu2016accelerate ; zhang2020new ; zhou2017unified . In particular, the Hölderian error bound with order $\delta\in(0,1]$, providing an upper bound for the distance of a point $x\in X$ to the solution set $\mathcal{X}^{*}$ in terms of the residual function, has been introduced for studying convergence rate of optimization methods; cf. johnstone2020faster . Special cases of this condition include the sharpness error bound with $\delta=1$, burke2002weak ; burke1993weak ; davis2018subgradient ; polyak1979sharp , and the quadratic growth condition with $\delta=\frac{1}{2}$, karimi2016linear ; zhang2020new . Considering these conditions, projected subgradient methods with carefully chosen step-sizes can achieve linear convergence under the extra weak convexity of the objective function; see, e.g., davis2018subgradient . As such, we will here investigate the convergence rate of the projected subgradient methods (with several different step-sizes) under Hölderian error bound and paraconvexity.
1.1 Contribution
Our primary goal is to extend the applicability of projected subgradient methods to a broader class of nonconvex, nonsmooth optimization problems, particularly focusing on paraconvex functions. We aim to provide comprehensive theoretical foundations to investigate convergence rates of the projected subgradient methods for such a class of functions and practical insights into this area. Our contribution has three folds, which we describe as follows:
  • (i)
    Characterization of the class of paraconvex functions. We first describe the class of paraconvex functions as a generalization to the class of weakly convex functions. Then, we verify that while preserving the concept of paraconvexity, the term $\min\{\lambda,1-\lambda\}$ can be removed from 1.2. Next, we establish the local Lipschitzness of such functions, ensuring the nonemptiness of their Clarke subdifferential. We particularly demonstrate a connection to a local version of $\nu$-paraconvex functions, known as $\nu$-weakly convex functions (see daniilidis2005filling ), coincide with $\nu$-paraconvex functions on compact convex sets. In addition, it is shown that the continuous midpoint $\nu$-paraconvexity implies $\nu$-paraconvexity, facilitating the assessment of paraconvexity. As a wider class than weakly convex function, they accept saddle point. As such, under Hölderian error bound, we identify a region around the optimal set containing no saddle points.
  • (ii)
    Convergence rate of projected subgradient methods with several step-sizes. We describe a generic framework for subgradient methods for paraconvex functions under the Hölderian error bound condition. In this setting, we study the convergence behavior of this subgradient method for several different step-sizes such as constant, nonsummable diminishing, square-summable but not summable, geometrically decaying, and Scaled Polyak’s step-sizes. We demonstrate recurrences to establish upper bounds for the sequence of iterations (i.e., $(\operatorname{dist}(x^{k},\mathcal{X}^{*}))_{k\in\mathbb{N}}$) and function values (i.e., $(f(x^{k})-f^{*})_{k\in\mathbb{N}}$) residuals. It is shown that: (i) for the projected subgradient method with constant step-size, $(\operatorname{dist}(x^{k},\mathcal{X}^{*}))_{k\in\mathbb{N}}$ is linearly convergent up to a specified tolerance; (ii) for the projected subgradient methods with nonsummable diminishing and square-summable but not summable step-sizes, the subsequential and global convergence are established, respectively; (iii) for the projected subgradient methods with geometrically decaying and Scaled Polyak’s step-sizes, $(\operatorname{dist}(x^{k},\mathcal{X}^{*}))_{k\in\mathbb{N}}$ is linearly convergent.
  • (iii)
    Applications to low-rank robust matrix recovery. We introduce a class of robust low-rank matrix recovery appearing in several applications such as robust matrix completion, image inpainting, matrix compression, and robust nonnegative matrix factorization. We apply the proposed projected subgradient method to these applications, and the results indicate the promising behavior of these algorithms. Surprisingly, the projected subgradient method with Scaled Polyak’s step-size performs much better than the algorithm with the original Polyak’s step-size. To the best of our knowledge, this is the first theoretical and numerical appearance of the projected subgradient method with Scaled Polyak’s step-size in the literature.
1.2 Paper organization
The remainder of the paper is organized as follows. Section 2 introduces the notations and reviews foundational concepts. In Section 3, we analyze the class of $\nu$-paraconvex functions and explore their fundamental properties. Section 4 focuses on projected subgradient methods for $\nu$-paraconvex functions satisfying the Hölderian error bound condition, deriving key inequalities. In Section 4.1, we investigate projected subgradient method with constant step-size. Section 4.2 discusses projected subgradient methods with diminishing step-size strategies. Section 4.3 examines projected subgradient methods using Scaled Polyak’s step-size. In Section 5, we present preliminary numerical results on robust low-rank matrix recovery validating our theoretical foundations. Finally, Section 6 delivers our concluding remarks.
2 Preliminaries
2.1 Notation
[…]
To derive the convergence rate bounds of our algorithms, we use the following fact from aragon2018accelerating .
Fact 2.1.
Let ${s_{k}}$ be a nonnegative real sequence and let $\alpha,\beta$ be some positive constants. Suppose that $s_{k}\to 0$ and that the sequence satisfies $s^{\alpha}_{k}\leq\beta(s_{k}-s_{k+1}),$ for all $k$ sufficiently large. Then, the following assertions hold:
  1. 1
    If $\alpha=0$, the sequence ${s_{k}}$ converges to 0 in a finite number of steps;
  2. 2
    If $\alpha\in(0,1]$, the sequence ${s_{k}}$ converges linearly to 0 with rate $1-\tfrac{1}{\beta}$;
  3. 3
    If $\alpha>1$, there exists $\eta>0$ such that $s_{k}\leq\eta k^{-\tfrac{1}{\alpha-1}}$, for all $k\in\mathbb{N}$ sufficiently large.
Let us conclude this subsection by addressing some concepts from variational analysis clarke1990optimization . Let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper function that is locally Lipschitz at $\bar{x}\in\operatorname{dom}(h)$. The generalized directional derivative of $h$ at $\bar{x}$ in the direction $d$ is given by $h^{\circ}(\bar{x};d):=\limsup_{\mathrel{{}\mathop{t\downarrow 0}\limits^{x\rightarrow\bar{x}}}}\frac{h(x+td)-h(x)}{t}$. Furthermore, the Clarke’s subdifferential of $h$ at $\bar{x}$ is defined as
$\partial h(\bar{x}):={\mathopen{}\left\{\zeta\in\mathcal{X}^{*}:\langle\zeta,d\rangle\leq h^{\circ}(\bar{x};d),\forall d\in\mathbb{R}^{n}\right\}\mathclose{}}.$
Moreover, given a nonempty set $S\subseteq\operatorname{dom}(h)$, a point $\bar{x}\in S$ is said to be a stationary point of the problem $\min_{x\in S}h(x)$ if $0\in\partial(h+\delta_{S})(\bar{x})$.
2.2 Hölderian error bound condition
An error bound is a crucial tool in optimization for achieving faster convergence rates in algorithms. In this study, we focus on a specific type of error bound known as the Hölderian error bound.
Definition 2.2.
A function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is said to admit a Hölderian error bound (HEB) of order $\delta\in(0,1]$ with constant $\mu>0$ on a nonempty set $S\subseteq\operatorname{dom}(h)$ if
$\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x;\mathcal{X}^{*})\leq h(x)-h^{*},\quad\forall x\in S,$ (2.1)
where $h^{*}:=\inf_{x\in S}h(x)$ and the set of minimizers $\mathcal{X}^{*}:=\{x\in S\,:\,h(x)=h^{*}\}$ is nonempty.
In the literature, this property is also known by other names, such as the Hölderian growth condition, Hölder growth bound, Hölder error bound, and global Hölderian error bound; see johnstone2020faster ; grimmer2023general . The HEB is instrumental in ensuring that projected subgradient methods with appropriate step-sizes achieve linear convergence, even for nonconvex problems davis2018subgradient . Two notable cases of HEB arise when $\delta=1$ and $\delta=\tfrac{1}{2}$. For $\delta=1$, the condition is known as the sharpness error bound or the weak sharpness condition burke1993weak ; polyak1979sharp , while for $\delta=\tfrac{1}{2}$, it is referred to as the quadratic growth condition karimi2016linear ; zhang2020new .
Beyond HEB (including the sharpness error bound and quadratic growth condition), other prominent properties have been studied, such as the Łojasiewicz inequality bolte2007lojasiewicz , the Kurdyka-Łojasiewicz (KL) inequality bolte2017error , and (Hölder) metric subregularity/regularity kruger2015error .
The Łojasiewicz inequality can be considered a version of HEB, where inequality (2.1) holds on each compact subset of an open set $S$. Furthermore, it has been established that HEB and the KL inequality are equivalent for convex, closed, and proper functions bolte2007lojasiewicz ; bolte2017error . On the other hand, Hölder metric subregularity represents a localized version of HEB. Additionally, (garrigos2023convergence, , Proposition 3.3) showed that HEB of order $\delta$ (referred to $\delta$-conditioned) implies Hölder metric subregularity (referred to $\delta$-metrically subregular) of $\partial f$ in the convex setting.
3 Fundamental properties of paraconvex functions and optimization
This section introduces the class of $\nu$-paraconvex functions and describes some of their key properties, which are necessary for the coming sections. Then, we consider a minimization of paraconvex functions and investigate a region in their domains that includes no saddle points.
3.1 Fundamental properties of paraconvex functions
Let us begin with the definition of $\nu$-paraconvexity.
Definition 3.1.
A function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is said to be $\nu$-paraconvex function on a convex set $S\subseteq\operatorname{dom}(h)$ for some $0<\nu\leq 1$ if there exists $\rho\geq 0$ such that for any $x,y\in S$ and $\lambda\in[0,1]$,
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+\rho\min\{\lambda,1-\lambda\}\|x-y\|^{1+\nu}.$ (3.1)
Remark 3.2.
To the best of our knowledge, Rolewicz rolewicz1979gamaparaconvex ; rolewicz1979paraconvex pioneered the study of paraconvexity, introducing this concept for the class of multi-valued mappings. In particular, for some $\gamma\geq 1$, a real-valued function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is said to be a $\gamma$-paraconvex function on a convex set $S\subseteq\operatorname{dom}(h)$, if there exists a constant $C\geq 0$ such that for any $x,y\in S$ and $\lambda\in[0,1]$,
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+C\|x-y\|^{\gamma}.$ (3.2)
In rolewicz2000alpha , the functions satisfying Definition 3.1 are referred to as strongly $\gamma$-paraconvex ($\gamma=1+\nu$). Notably, Proposition 3.3 establishs $\gamma$-paraconvexity and strong $\gamma$-paraconvexity are equivalent when $1<\gamma\leq 2$ ($0<\nu\leq 1$). Furthermore, these properties reduce to convexity when $\gamma>2$ ($\nu>1$) due to (rolewicz2000alpha, , Proposition 7). In addition, Jourani (jourani1996subdifferentiability, , Corollary 7.1) demonstrated that a function $h$ is $\gamma$-paraconvex on $S$ with $\gamma>1$ if and only if there exists a constant $C\geq 0$ such that for all $x,y\in S$ and $0\leq\lambda\leq 1$,
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+C\lambda(1-\lambda)\|x-y\|^{\gamma}.$ (3.3)
Here, we adopt the definition of $\nu$-paraconvexity as given in expression 3.1 of Definition 3.1, where $0<\nu\leq 1$ ($1<\gamma\leq 2$). This formulation is motivated by the fact that the term $\min\{\lambda,(1-\lambda)\}$ plays a key role in 3.1, with equality holding at $\lambda=0$ and $\lambda=1$. Furthermore, several important properties and characterizations, discussed later in this section, hold when $\gamma>1$. ∎
The next result shows that the concept of $\nu$-paraconvexity remains unchanged if the term $\min\{\lambda,1-\lambda\}$ is omitted from the inequality in 3.1. Without proof, this result was presented in rolewicz1979gamaparaconvex ; jourani1996subdifferentiability .
Proposition 3.3.
A function $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is a $\nu$-paraconvex function on a convex set $S$ if and only if there exists a constant $C\geq 0$ such that for any $x,y\in S$ and $\lambda\in[0,1]$,
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+C\|x-y\|^{1+\nu}.$ (3.4)
Proof.
The necessity is clear from the elementary fact that $\min\{\lambda,1-\lambda\}\leq 1$. For the sufficiency, assume that the expression 3.4 holds. Fixing any $x,y\in S$ and $\lambda\in[0,1]$, we will verify 3.1. We can suppose, without loss of generality, that $0<\lambda\leq\tfrac{1}{2}$. So, taking the function $\varphi:[0,1]\rightarrow\mathbb{R}$ defined as $\varphi(t):=h(z_{t})-th(x)-(1-t)h(y)$ in which $z_{t}=tx+(1-t)y$, the argument reduces to showing that there exists some $\rho\geq 0$ such that $\varphi(\lambda)\leq\rho\lambda\|x-y\|^{1+\nu}$. It is evident that $\varphi(0)=\varphi(1)=0$, $\varphi(\tfrac{1}{2})\leq C\|x-y\|^{1+\nu}$, by 3.4, and there exists $\kappa\in\mathbb{N}$ such that $\lambda\in[2^{-\kappa},2^{-\kappa+1})$ having the expression $\lambda=\mu 2^{-\kappa}+(1-\mu)2^{-\kappa+1}$ for some $\mu\in(0,1]$. Applying induction, we show
$\varphi(\tfrac{1}{2^{k}})\leq\tfrac{2^{1+\nu}}{2^{k}}C\|x-y\|^{1+\nu}\sum_{i=1}^{k}(\tfrac{1}{2^{\nu}})^{i}.$ (3.5)
Then, it follows that
$\varphi(\lambda)$ $=\varphi(\mu\tfrac{1}{2^{\kappa}}+(1-\mu)\tfrac{1}{2^{\kappa-1}})=h(\mu z_{2^{-\kappa}}+(1-\mu)z_{2^{1-\kappa}})-\lambda h(x)-(1-\lambda)h(y)$
$\leq\mu h(z_{2^{-\kappa}})+(1-\mu)h(z_{2^{1-\kappa}})+C\|z_{2^{-\kappa}}-z_{2^{1-\kappa}}\|^{1+\nu}-\lambda h(x)-(1-\lambda)h(y)$
$=\mu\varphi(\tfrac{1}{2^{\kappa}})+(1-\mu)\varphi(\tfrac{1}{2^{\kappa-1}})+\tfrac{C}{2^{\kappa(1+\nu)}}\|x-y\|^{1+\nu}\leq\bigg{(}\mu\tfrac{2^{1+\nu}}{2^{\kappa}}\sum_{i=1}^{\kappa}(\tfrac{1}{2^{\nu}})^{i}+(1-\mu)\tfrac{2^{1+\nu}}{2^{\kappa-1}}\sum_{i=1}^{\kappa-1}(\tfrac{1}{2^{\nu}})^{i}+\tfrac{1}{2^{\kappa(1+\nu)}}\bigg{)}C\|x-y\|^{1+\nu}$
$\leq\bigg{(}\mu 2^{1+\nu}\sum_{i=1}^{\kappa}(\tfrac{1}{2^{\nu}})^{i}+(1-\mu)2^{2+\nu}\sum_{i=1}^{\kappa-1}(\tfrac{1}{2^{\nu}})^{i}+\tfrac{1}{2^{\kappa\nu}}\bigg{)}\lambda C\|x-y\|^{1+\nu}\leq\rho\lambda\|x-y\|^{1+\nu},$
where $\rho:=\tfrac{42^{\nu}}{1-2^{\nu}}C$, as desired. To show 3.5, we assume that it holds for $k\in\mathbb{N}$. Invoking 3.4, for each $t\in[0,1]$, we come to
$\varphi(\tfrac{t}{2})=h{\mathopen{}\left(\tfrac{z_{t}+y}{2}\right)\mathclose{}}-\tfrac{t}{2}h(x)-(1-\tfrac{t}{2})h(y)\leq\tfrac{1}{2}h(z_{t})+\tfrac{1}{2}h(y)+Ct^{1+\nu}\|x-y\|^{1+\nu}-\tfrac{t}{2}h(x)-(1-\tfrac{t}{2})h(y)=\tfrac{1}{2}\varphi(t)+Ct^{1+\nu}\|x-y\|^{1+\nu}.$
Employing the above inequality and inductive assumption, it holds that
$\varphi(\tfrac{1}{2^{k+1}})=\varphi\Big{(}\tfrac{\tfrac{1}{2^{k}}}{2}\Big{)}\leq\tfrac{1}{2}\varphi(\tfrac{1}{2^{k}})+\frac{C}{2^{k(1+\nu)}}\|x-y\|^{1+\nu}\leq\bigg{(}\tfrac{2^{1+\nu}}{2^{k+1}}\sum_{i=1}^{k}(\tfrac{1}{2^{\nu}})^{i}+\frac{1}{2^{k(1+\nu)}}\bigg{)}C\|x-y\|^{1+\nu}=\tfrac{2^{1+\nu}}{2^{k+1}}C\|x-y\|^{1+\nu}\sum_{i=1}^{k+1}(\tfrac{1}{2^{\nu}})^{i},$
adjusting our desired result. ∎
The class of $\nu$-paraconvex functions is significant in various applications, though they may not always be straightforward to identify. In the following, we provide some notable examples of this class of functions in different contexts.
Example 3.4 (Known paraconvex functions).
We next mention several classes of paraconvex functions.
  • (a)
    One notable example of $\nu$-paraconvex functions lies in the class of composite optimization problems:
    $\min_{x\in S}h(x):=\varphi(g(x)),$ (3.6)
    where $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is a convex and locally Lipschitz function with constant $L_{0}$, and the Jacobian of $g:\mathbb{R}^{m}\to\mathbb{R}^{n}$, $J(x)$, is Hölderian continuous of order $\nu\in(0,1]$ with constant $L_{1}$ on the convex set $S$:
    $\|J(x)-J(y)\|\leq L\|x-y\|^{\nu},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall x,y\in S.$
    It can be shown that the function $h=\varphi\circ g$ is $\nu$-paraconvex with constant $\tfrac{2L_{0}L_{1}}{\nu+1}$ on $S$; see Proposition 3.14.
  • (b)
    Any smooth function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ whose gradient is Hölderian continuous of order $\nu\in(0,1]$ with constant $L$ on a convex set $S$, is $\nu$-paraconvex function with constant $\tfrac{2L}{\nu+1}$ on $S$.
  • (c)
    In robust principal component analysis (Robust PCA), the objective is to detect sparse corruptions in a low-rank matrix. Mathematically, given an $m\times n$ matrix $A$, the task is to decompose it into $A=UV^{T}+S$ where $U$ and $V$ are low rank matrices and $S\in\mathbb{R}^{m\times n}$ is a sparse matrix. This problem can be formulated as:
    $\min\|UV^{T}-(A-S)\|_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ U\in\mathbb{R}^{m\times r},V\in\mathbb{R}^{n\times r},$
    where $r$ denotes the desired rank. The objective function of this problem is a $1-$paraconvex function and satisfies HEB of order $\delta=1$; see drusvyatskiy2017proximal .
  • (d)
    As another example, consider phase retrieval as a widely encountered computational problem. This problem appears in real-world applications in various fields like speech processing, and imaging. Given the vectors $a_{i}\in\mathbb{R}^{n}$, $i=1,2,\ldots,m,$ and $b\in\mathbb{R}^{m}$, the real-valued version of the problem concentrates on finding a solution $x$ which satisfies the following magnitude conditions:
    $\langle a_{i},x\rangle^{2}\approx b_{i},\leavevmode\nobreak\ \forall i=1,2,\ldots,m.$
    In this situation, the problem can be formulated as davis2020nonsmooth ; duchi2019solvingg ; eldar2014phase :
    $\min_{x}\tfrac{1}{m}\sum_{i=1}^{m}|\langle a_{i},x\rangle^{2}-b_{i}|=\min_{x}\tfrac{1}{m}\|(Ax)^{2}-b\|_{1},$
    where $A$ is a $m\times n$ matrix whose $i-$th row is $a^{T}_{i}$ and $(Ax)^{2}$ is a vector in $\mathbb{R}^{m}$ whose $i-$th component is $\langle a_{i},x\rangle^{2}$. The objective function of this problem is an example of composite problem 3.6. Invoking (duchi2019solvingg, , Corollary 3.2 and Proposition 3), it can be seen that this function is $1-$paraconvex and holds HEB of order $1$ under some reasonable statistical conditions by considering a suitable noise model in the measurements.
Remark 3.5.
Rolewicz rolewicz1979gamaparaconvex ; rolewicz1979paraconvex demonstrated that $1-$paraconvex functions (when $\nu=1$) can be represented as the difference of two convex functions. Specifically, a function $h$ is $1-$paraconvex if and only if there exists a nonnegative constant $C$ such that $h(x)+C\|x\|^{2}$ is a convex function. In the literature, these functions are also referred to as weakly convex functions nurminskii1973quasigradient . Furthermore, they are equivalent to $LC^{2}$ functions, as introduced by Rockafellar rockafellar1981favorable . For further details on the role of these functions in optimization, see rockafellar1981favorable ; hiriart1985generalized ; penot1996favorable ; daniilidis2005filling .
For the case that $10<\nu<$ 1, Rolewicz (rolewicz1979gamaparaconvex, , Example 1) presented an example of a $\nu$-paraconvex function which it cannot be expressed as the difference of two convex functions. However, we establish that the convexity of $h(x)+C\|x\|^{\nu+1}$ for some $10<\nu<$ 1 and $C\geq 0$ implies the $\nu$-paraconvexity of $h$ (see Proposition 3.6). Moreover, in Example 3.8, we provide an example that highlights this result, emphasizing the existence of $\nu$-paraconvex functions with $10<\nu<$ 1 that are not $1-$paraconvex (weakly convex). ∎
Next, we show that the convexity of the function $x\mapsto h(x)+C\|x\|^{\nu+1}$ is sufficient for the $\nu$-paraconvexity of $h$.
Proposition 3.6.
For a given function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$, if there exist $10<\nu<$ 1 and $C\geq 0$ such that $x\mapsto h(x)+C\|x\|^{\nu+1}$ is convex, then, $h$ is $\nu$-paraconvex on $\operatorname{int}\operatorname{dom}(h)$.
Proof.
Let $h(\cdot)+C\|\cdot\|^{\nu+1}$ be a convex function, and let $x,y\in\operatorname{int}\operatorname{dom}(h)$ and $10<\lambda<$ 1, i.e.,
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+C{\mathopen{}\left(\lambda\|x\|^{\nu+1}+(1-\lambda)\|y\|^{\nu+1}-\|\lambda x+(1-\lambda)y\|^{\nu+1}\right)\mathclose{}}.$ (3.7)
Applying the inequality $(a+b)^{p}\leq a^{p}+b^{p}$ for $a,b\geq 0$ and $0<p\leq 1$ (with $p=\frac{\nu+1}{2}$), along with the concavity of the function $t^{\tfrac{\nu+1}{2}}$ on $[0,+\infty)$, we obtain
$\|\lambda x+(1-\lambda)y\|^{\nu+1}+\sqrt{\lambda^{\nu+1}(1-\lambda)^{\nu+1}}\|x-y\|^{\nu+1}$ $\geq{\mathopen{}\left(\|\lambda x+(1-\lambda)y\|^{2}+\lambda(1-\lambda)\|x-y\|^{2}\right)\mathclose{}}^{\tfrac{\nu+1}{2}}$
$={\mathopen{}\left(\lambda\|x\|^{2}+(1-\lambda)\|y\|^{2}\right)\mathclose{}}^{\tfrac{\nu+1}{2}}\geq\lambda\|x\|^{\nu+1}+(1-\lambda)\|y\|^{\nu+1}.$
Together with 3.7, this implies
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+C\sqrt{\lambda^{\nu+1}(1-\lambda)^{\nu+1}}\|x-y\|^{\nu+1},$
which ensures the $\nu$-paraconvexity of $h$ as stated in Proposition 3.3. ∎
Corollary 3.7.
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a convex function on a convex set $S\subseteq\operatorname{dom}(h)$. Then, for each $\nu\in(0,1]$ and $C\geq 0$, the function $x\mapsto h(x)-C\|x\|^{\nu+1}$ is $\nu$-paraconvex on $\operatorname{int}S$.
Example 3.8.
Let us consider the function $h:\mathbb{R}\rightarrow\mathbb{R}$ given by
$h(x)={\mathopen{}\left\{\begin{array}[]{ll}1-|x|^{1.5},&-1\leq x\leq 1,\\ x^{2}-1,&o.w.\end{array}\right.\mathclose{}}$
The function $x\mapsto h(x)+|x|^{1.5}$ is evidently convex, which implies that $h$ is a $0.5$-paraconvex function. Furthermore, it can be demonstrated that for any $\rho\geq 0$, the function $x\mapsto h(x)+\rho\|x\|^{2}$ is not convex and possesses some saddle points. This indicates that the original function $h$ is not $1$-paraconvex (i.e., weakly convex). Nonetheless, the function $x\mapsto h(x)+\rho|x|^{2}$ also qualifies as a $0.5$-paraconvex function.
Figure 1: Comparisons of the function $h(x)$ and its modified versions demonstrating $\nu$-paraconvexity. The left plot shows $h(x)$ (in black), $h(x)+|x|^{1.5}$ (in green), highlighting the convexity of $h(x)+|x|^{1.5}$ which implies $h$ is $0.5$-paraconvex. The right plot shows $h(x)$ alongside $h(x)+|x|^{2}$ (in purple), illustrating that $h$ is not $1$-paraconvex.
Remark 3.9.
A further extension of convexity in the nonconvex setting is $\nu$-weak convexity daniilidis2005filling . Given a nonempty open set $\mathcal{U}\subseteq\mathbb{R}^{n}$ and $0<\nu\leq 1$, a locally Lipschitz function $\varphi:\mathcal{U}\rightarrow\mathbb{R}$ is said to be $\nu$-weakly convex, if for any $x_{0}\in\mathcal{U}$, there exist $\delta>0$ and $C\geq 0$ such that for all $x,y\in\mathbb{B}(x_{0};\delta)$ and $\lambda\in[0,1]$,
$\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)+C\lambda(1-\lambda)\|x-y\|^{1+\nu}.$
This concept is a local version of $\nu$-paraconvexity. Hence, $\nu$-paraconvexity implies $\nu$-weak convexity. Furthermore, Theorem 3.11 demonstrates that the converse holds on compact convex sets. ∎
Lemma 3.10.
Let $\mathcal{I}$ be an interval in $\mathbb{R}$ and $a,b\in I$ such that $a<b$. Let $\varphi:\mathcal{I}\rightarrow\mathbb{R}$ be a $\nu$-paraconvex function on $\mathcal{I}\cap[a,+\infty)$ and $\mathcal{I}\cap(-\infty,b]$ for some $\nu\in(0,1]$ with constant $C\geq 0$ (comes from 3.3). Then, $\varphi$ is a $\nu$-paraconvex function on $\mathcal{I}$ with constant $3C$.
Proof.
If $a=\inf\mathcal{I}$ or $b=\sup\mathcal{I}$, then there is nothing to prove. Therefore, without loss of generality, assume that $a,b\in\operatorname{int}\mathcal{I}$. Let $x,y\in\mathcal{I}$ with $x<y$, and let $10<\lambda<$ 1. Set $z=\lambda x+(1-\lambda)y$. We aim at showing that
$\varphi(z)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)+3C\lambda(1-\lambda)(y-x)^{\nu+1}.$ (3.8)
If $a<x$ or $y<b$, the inequality 3.8 trivially holds due to the $\nu$-paraconvexity of $h$ on $\mathcal{I}\cap[a,+\infty)$ and $\mathcal{I}\cap(-\infty,b]$. Now, assume that $x<a<b<y$. Three possible cases arise: (i) $z\in(x,a)$; (ii) $z\in[a,b]$; (iii) $z\in(b,y)$.
In Case (i), $z\in(x,a)$, there exist $r,s,t\in(0,1)$ such that $z=ra+(1-r)x$, $a=sb+(1-s)x$, and $b=ty+(1-t)a$. By substituting these relations, we derive: $a=sty+s(1-t)a+(1-s)x$ and $b=ty+(1-t)sb+(1-t)(1-s)x,$ leading to
$a=\tfrac{st}{1-s(1-t)}y+\tfrac{1-s}{1-s(1-t)}x,\quad b=\tfrac{t}{1-s(1-t)}y+\tfrac{(1-t)(1-s)}{1-s(1-t)}x,\quad z={\mathopen{}\left((1-r)+\tfrac{r(1-s)}{1-s(1-t)}\right)\mathclose{}}x+\tfrac{rst}{1-s(1-t)}y.$ (3.9)
Thus, $\lambda=(1-r)+\tfrac{r(1-s)}{1-s(1-t)}$. Furthermore, the $\nu$-paraconvexity of $\varphi$ on $\mathcal{I}\cap[a,+\infty)$ and $\mathcal{I}\cap(-\infty,b]$ verifies that
$\,\,\varphi(b)\leq t\varphi(y)+(1-t)\varphi(a)+Ct(1-t)(y-a)^{\nu+1},$ (3.10)
$\varphi(a)\leq s\varphi(b)+(1-s)\varphi(x)+Cs(1-s)(b-x)^{\nu+1},$ (3.11)
$\,\varphi(z)\leq r\varphi(a)+(1-r)\varphi(x)+Cr(1-r)(a-x)^{\nu+1}.$ (3.12)
Combining 3.10 and 3.11 ensures
$\varphi(a)\leq s{\mathopen{}\left(t\varphi(y)+(1-t)\varphi(a)+Ct(1-t)(y-a)^{\nu+1}\right)\mathclose{}}+(1-s)\varphi(x)+Cs(1-s)(b-x)^{\nu+1},$
and consequently,
$\varphi(a)\leq\tfrac{st}{1-s(1-t)}\varphi(y)+\tfrac{1-s}{1-s(1-t)}\varphi(x)+C{\mathopen{}\left(\tfrac{st(1-t)}{1-s(1-t)}(y-a)^{\nu+1}+\tfrac{s(1-s)}{1-s(1-t)}(b-x)^{\nu+1}\right)\mathclose{}}.$
Now, the inequality 3.8 is confirmed via 3.12, i.e.,
$\varphi(z)$ $\leq{\mathopen{}\left((1-r)+\tfrac{r(1-s)}{1-s(1-t)}\right)\mathclose{}}\varphi(x)+\tfrac{rst}{1-s(1-t)}\varphi(y)+C{\mathopen{}\left(\tfrac{rst(1-t)}{1-s(1-t)}(y-a)^{\nu+1}+\tfrac{rs(1-s)}{1-s(1-t)}(b-x)^{\nu+1}+r(1-r)(a-x)^{\nu+1}\right)\mathclose{}}$
$\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)+3C\lambda(1-\lambda)(y-x)^{\nu+1}.$
To show Case (ii), we assume that $z\neq a$. There exist $r,s\in(0,1)$ such that $z=ra+(1-r)y$ and $a=sx+(1-s)z$ implying $z=\tfrac{rs}{1-r(1-s)}x+\tfrac{1-r}{1-r(1-s)}y.$ So, $\lambda=\tfrac{rs}{1-r(1-s)}$. Moreover, relying again on $\nu$-paraconvexity of $\varphi$ we deduce
$\varphi(z)\leq r\varphi(a)+(1-r)\varphi(y)+Cr(1-r)(y-a)^{\nu+1}\leq r{\mathopen{}\left(s\varphi(x)+(1-s)\varphi(z)+Cs(1-s)(z-x)^{\nu+1}\right)\mathclose{}}+(1-r)\varphi(y)+Cr(1-r)(y-a)^{\nu+1},$
which comes out as 3.8. In the case where $z=a$, we have $z\neq b$, and following a similar argument yields the desired. The proof of Case (iii) is similar to Cases (ii)-(iii) and is therefore omitted. ∎
Our next result indicates that $\nu$-weak convexity leads to $\nu$-paraconvexity on the compact convex sets.
Theorem 3.11 ($\nu$-weak convexity and $\nu$-paraconvexity).
Let $\mathcal{U}\subseteq\mathbb{R}^{n}$ be an open set and $S\subseteq\mathcal{U}$ be a nonempty convex and compact set. Let $h:\mathcal{U}\rightarrow\mathbb{R}$ be a $\nu$-weakly convex function for some $\nu\in(0,1]$. Then, $h$ is a $\nu$-paraconvex function on $S$.
Proof.
Let us begin by showing that the inequality of $\nu$-weak convexity holds uniformly on $S$, with a fixed constant $C\geq 0$ independent of $x\in S$. The $\nu$-weak convexity of $h$ implies that for each point $x\in S$, there exist constants $C_{x}\geq 0$ and $\delta_{x}>0$ such that for all $x_{0},y_{0}\in\mathbb{B}(x;\delta_{x})$ and $\lambda\in[0,1]$,
$h(\lambda x_{0}+(1-\lambda)y_{0})\leq\lambda h(x_{0})+(1-\lambda)h(y_{0})+C_{x}\lambda(1-\lambda)\|x_{0}-y_{0}\|^{1+\nu}.$ (3.13)
The collection of open balls, $\{\mathbb{B}(x;\delta_{x})\}_{x\in S}$ covers $S$, and by the compactness principle a finite subcollection of them covers $S$. Thus, there exist points $x_{i}\in S$ and radii $\delta_{i}>0$, $i=1,2,\ldots,m,$ such that $S\subseteq\bigcup_{i=1}^{m}\mathbb{B}(x_{i};\delta_{i})$. Define $\overline{C}:=\max\{C_{i}:\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,m\}$. Then, for each $x\in S$, there exists some $\delta_{x}>0$ such that $\mathbb{B}(x;\delta_{x})\subseteq\mathbb{B}(x_{i};\delta_{i})$, namely for an index $i$ such that $x\in\mathbb{B}(x_{i};\delta_{i})$, and 3.13 holds on $\mathbb{B}(x;\delta_{x})$ with the fixed constant $\overline{C}$.
Let $x,y\in S$ and $\mu\in[0,1]$ be arbitrary and constant hereafter. Let us consider the function $\varphi:[0,1]\to\mathbb{R}$ defined as $\varphi(t):=h(z_{t})$, where $z_{t}=y+t(x-y)$ and $t\in[0,1]$. We aim at showing that $\varphi$ is $\nu$-paraconvex on $[0,1]$ with constant $9\overline{C}\|y-x\|^{\nu+1}$. Once this is established, we have
$h(z_{\mu})=\varphi(\mu)\leq\mu\varphi(1)+(1-\mu)\varphi(0)+9\overline{C}\|y-x\|^{\nu+1}\mu(1-\mu)|1-0|^{\nu+1}=\mu h(x)+(1-\mu)h(y)+9\overline{C}\mu(1-\mu)\|y-x\|^{\nu+1},$
as our desired result. Let $\hat{t}\in(0,1)$. Inasmuch as $S$ is convex, it follows that $z_{\hat{t}}\in S$. So, there exists $\delta>0$ such that for all $x_{0},y_{0}\in\mathbb{B}(z_{\hat{t}};\delta)$ and $\lambda\in[0,1]$, the inequality 3.13 is satisfied with constant $\overline{C}$. Let us consider $\epsilon:={\tfrac{1}{2}}\min\{1-\hat{t},\hat{t},\tfrac{\delta}{\|y-x\|}\}$. Suppose $t_{1},t_{2}\in[\hat{t}-\epsilon,\hat{t}+\epsilon]$ and $\lambda\in[0,1]$. Then, $z_{t_{1}},z_{t_{2}}\in\mathbb{B}(z_{\hat{t}};\delta)$ and therefore
$h(\lambda z_{t_{1}}+(1-\lambda)z_{t_{2}})\leq\lambda h(z_{t_{1}})+(1-\lambda)h(z_{t_{2}})+\overline{C}\lambda(1-\lambda)\|z_{t_{1}}-z_{t_{2}}\|^{\nu+1},$
implying
$\varphi(\lambda t_{1}+(1-\lambda)t_{2})\leq\lambda\varphi(t_{1})+(1-\lambda)\varphi(t_{2})+\overline{C}\|y-x\|^{\nu+1}\lambda(1-\lambda)|t_{1}-t_{2}|^{\nu+1}.$
Hence, $\varphi$ is $\nu$-paraconvex on $[\hat{t}-\epsilon,\hat{t}+\epsilon]$ with constant $\overline{C}\|y-x\|^{\nu+1}$. Now, let us define
$A:={\mathopen{}\left\{t\in[0,1]:\leavevmode\nobreak\ \varphi\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \nu-\text{paraconvex on}\leavevmode\nobreak\ [\hat{t}-\epsilon,t]\leavevmode\nobreak\ \text{with constant}\leavevmode\nobreak\ 3\overline{C}\|y-x\|^{\nu+1}\right\}\mathclose{}},$
and consider $a:=\sup A$. Clearly, $\hat{t}+\epsilon\in A$. So, the supremum is well-defined and $0<a\leq 1$. It is easy to see that $a\in A$ using continuity of $h$ (see Proposition 3.12). We claim that $a=1$. Otherwise, if $a\in(0,1)$, the argument applied for $\hat{t}$ can similarly be repeated with $a$ to conclude that $\varphi$ is $\nu$-paraconvex on $[a-\bar{\epsilon},a+\bar{\epsilon}]\subseteq(0,1)$ for some $\bar{\epsilon}>0$ with constant $\overline{C}\|y-x\|^{\nu+1}$. In this case, the function $\varphi$ is $\nu$-paraconvex on $[\hat{t}-\epsilon,a]$ and $[a-\bar{\epsilon},a+\bar{\epsilon}]$ with constant $\overline{C}\|y-x\|^{\nu+1}$, and so on $[\hat{t}-\epsilon,a+\bar{\epsilon}]$ with constant $3\overline{C}\|y-x\|^{\nu+1}$, by Lemma 3.10. This contradicts with $a=\sup A$. Thus, $a=1$. So, $\varphi$ is $\nu$-paraconvex on $[\hat{t}-\epsilon,1]$ with constant $3\overline{C}\|y-x\|^{\nu+1}$. Similarly, it can be shown that $\varphi$ is $\nu$-paraconvex on $[0,\hat{t}+\epsilon]$ with constant $3\overline{C}\|y-x\|^{\nu+1}$. Hence, $\varphi$ is $\nu$-paraconvex on $[0,1]$ with constant $9\overline{C}\|y-x\|^{\nu+1}$, due to Lemma 3.10, and the proof is completed. ∎
In what follows, we investigate several properties and characterizations of the class of $\nu$-paraconvex functions. First, similar to the convex case, we observe that $\nu$-paraconvex functions exhibit local Lipschitz continuity. Although this property was proven in jourani1996subdifferentiability ; jourani1996open under the assumptions of local boundedness or continuity, we establish here that $\nu$-paraconvex functions are locally bounded, thereby directly satisfying the local Lipschitz continuity property.
Proposition 3.12 (Locally bounded and Lipschitz properties).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a $\nu$-paraconevex function for some $\nu\in(0,1]$ and constant $\rho\geq 0$ on a convex set $S\subseteq\mathbb{R}^{n}$. Then the following assertions hold:
  • (a)
    The function $h$ is locally bounded on $\operatorname{int}S$.
  • (b)
    The function $h$ is locally Lipschitz on $\operatorname{int}S$.
Proof.
(a) Let $\bar{x}\in\operatorname{int}S$. There exists $\epsilon>0$ such that $P:=\{x\in\mathbb{R}^{n}:\|x-\bar{x}\|_{\infty}\leq\epsilon\}\subseteq S.$ Clearly, polyhedral $P$ has $2^{n}$ vertexes. Assume that $E$ is the set of all vertexes of the polyhedral $P$ and $M:=\max_{x\in E}h(x).$ Let $x\in P$. Then, thank to the Minkowski-Weyl representation (bertsekas2003convex, , Proposition 3.2.2) and extension of the Caratheodory’s theorem (bertsekas2003convex, , Theorem 1.22), there exist the scalars $\lambda_{j}\in[0,1]$ and the vertexes $x^{j}\in E$, $j=1,2\ldots,n+1$, such that $x=\sum_{j=1}^{n+1}\lambda_{j}x^{j}$ and $\sum_{j=1}^{n+1}\lambda_{j}=1.$ Using approximate Jensen’s inequality (van2022variational, , Lemma 2.1), we deduce
$h(x)\leq\sum_{j=1}^{n+1}\lambda_{j}h(x^{j})+\rho\sum_{j=1}^{n+1}\lambda_{j}(1-\lambda_{j})\max_{1\leq k\leq n+1}\|x^{k}-x^{j}\|^{1+\nu}\leq M+\frac{n+1}{4}\sqrt{n}^{1+\nu}2^{1+\nu}\epsilon^{1+\nu}\rho.$
Now, assume that $x\in cl\mathbb{B}(\bar{x};\epsilon)\subseteq P$. Set, $\bar{\lambda}:=\frac{\epsilon}{\epsilon+\|x-\bar{x}\|}$. Then,
$h(\bar{x})\leq\bar{\lambda}h(x)+(1-\bar{\lambda})h{\mathopen{}\left(\tfrac{1}{1-\bar{\lambda}}\bar{x}-\tfrac{\bar{\lambda}}{1-\bar{\lambda}}x\right)\mathclose{}}+\rho\min\{\bar{\lambda},1-\bar{\lambda}\}{\mathopen{}\left\|x-\tfrac{1}{1-\bar{\lambda}}\bar{x}+\tfrac{\bar{\lambda}}{1-\bar{\lambda}}x\right\|\mathclose{}}^{1+\nu},$
i.e.,
$h(x)$ $\geq\tfrac{1}{\bar{\lambda}}h(\bar{x})-\tfrac{1-\bar{\lambda}}{\bar{\lambda}}h{\mathopen{}\left(\tfrac{1}{1-\bar{\lambda}}\bar{x}-\tfrac{\bar{\lambda}}{1-\bar{\lambda}}x\right)\mathclose{}}-\tfrac{\rho\min\{\bar{\lambda},(1-\bar{\lambda})\}}{\bar{\lambda}}{\mathopen{}\left(\epsilon+\|x-\bar{x}\|\right)\mathclose{}}^{1+\nu}\geq h(\bar{x})-M-{\mathopen{}\left(\tfrac{n+1}{4}\sqrt{n}^{1+\nu}-1\right)\mathclose{}}2^{1+\nu}\epsilon^{1+\nu}\rho.$
Setting
$\overline{M}:=\max{\mathopen{}\left\{M+\tfrac{n+1}{4}\sqrt{n}^{1+\nu}2^{1+\nu}\epsilon^{1+\nu}\rho,{\mathopen{}\left|h(\bar{x})-M-{\mathopen{}\left(\tfrac{n+1}{4}\sqrt{n}^{1+\nu}-1\right)\mathclose{}}2^{1+\nu}\epsilon^{1+\nu}\rho\right|\mathclose{}}\right\}\mathclose{}},$
we come to $|h(x)|\leq\overline{M}$ for all $x\in cl\,\mathbb{B}(\bar{x};\epsilon)$, i.e., $h$ is locally bounded on $\operatorname{int}S$. Assertion (b) follows from Assertion (a) in combination with (jourani1996open, , Lemma 2.5) and (jourani1996subdifferentiability, , Proposition 2.2), adjusting our desired result. ∎
We next present some characterizations of the $\nu$-paraconvexity that are the straightforward consequences of the definition. Some of these results have been partially explored in the recent work by jourani1996subdifferentiability , which we include here for completeness.
Proposition 3.13 (Paraconvexity characterization).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a locally Lipschitz function defined on a nonempty convex set $S\subseteq\operatorname{dom}(h)$. Then, the following statements hold:
  • (a)
    If $h$ is a $\nu$-paraconvex function on $S$ for some $\nu\in(0,1]$ with constant $\rho\geq 0$, then for any $x,y\in S$ and any $\zeta\in\partial(h+\delta_{S})(x)$,
    $h(y)\geq h(x)+\langle{}\zeta{},{}y-x{}\rangle-\rho\|y-x\|^{\nu+1}.$ (3.14)
    Conversely, if inequality 3.14 holds, then $f$ is $\nu$-paraconvex function on $S$ with constant $2\rho$.
  • (b)
    The function $h$ is a $\nu$-paraconvex function on $S$ for some $\nu\in(0,1]$ if and only if $\partial h$ is $\nu$-paramonotone on $S$, i.e., there exists some $C\geq 0$ such that for any $x,y\in S$, any $\zeta\in\partial(h+\delta_{S})(x)$, and any $\eta\in\partial(h+\delta_{S})(y)$,
    $\langle{}\eta-\zeta{},{}y-x{}\rangle\geq-C\|x-y\|^{\nu+1}.$ (3.15)
  • (c)
    If $h$ is a twice continuously differentiable function on $S$ and for some $\nu\in(0,1]$ and $\rho\geq 0$, the matrix
    $\nabla^{2}h(x)+\rho\tfrac{1+\nu}{\|x\|^{3-\nu}}{\mathopen{}\left(\|x\|^{2}I-(1-\nu)xx^{T}\right)\mathclose{}},$
    is positive semidefinite for each $x\in S$ with $x\neq 0$, then $h$ is a $\nu$-paraconvex function on $S$.
Proof.
(a) For the first implication, see (jourani1996subdifferentiability, , Theorem 3.1). Let $x,y\in S$ and $10<\lambda<$ 1 be arbitrary to verify the converse. Let us consider $\zeta\in\partial(h+\delta_{S})(\lambda x+(1-\lambda)y)$. On the basis of 3.14, we obtain
$h(x)$ $\geq h(\lambda x+(1-\lambda)y)-(1-\lambda)\langle{}\zeta{},{}y-x{}\rangle-\rho(1-\lambda)^{\nu+1}\|y-x\|^{\nu+1},$
$\leavevmode\nobreak\ h(y)\geq h(\lambda x+(1-\lambda)y)+\lambda\langle{}\zeta{},{}y-x{}\rangle-\rho\lambda^{\nu+1}\|y-x\|^{\nu+1},$
which, by combining these inequalities, gives us
$\begin{array}[]{ll}\lambda h(x)+(1-\lambda)h(y)&\geq h(\lambda x+(1-\lambda)y)-\rho(\lambda(1-\lambda)^{\nu+1}+(1-\lambda)\lambda^{\nu+1})\,\|y-x\|^{\nu+1},\\[5.69054pt] &\geq h(\lambda x+(1-\lambda)y)-2\rho\min\{\lambda,(1-\lambda)\}\,\|y-x\|^{\nu+1}.\end{array}$
(b) See (jourani1996subdifferentiability, , Coeollary 7.1). (c) The proof of this assertion follows directly from Proposition 3.6. ∎
Now, applying this characterization, we can show the $\nu$-paraconvexity of a composite function.
Proposition 3.14 (Composite paraconvex function).
Let $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ be a convex and locally Lipschitz function with constant $L_{0}$. Assume that $g:\mathbb{R}^{m}\to\mathbb{R}^{n}$ has Hölderian continuous Jacobian of order $\nu\in(0,1]$ with constant $L_{1}$ on $S$. Then, the composite function $h:=\varphi\circ g$ is $\nu$-paraconvex with constant $\rho=\tfrac{2L_{0}L_{1}}{\nu+1}$ on $S$.
Proof.
The Hölderian continuity of Jacobian of $g$ ensures
$\|g(y)-g(x)-J(x)(y-x)\|\leq\tfrac{L_{1}}{\nu+1}\|y-x\|^{\nu+1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall x,y\in S,$
as stated in (yashtini2016global, , Lamma 1). Let $x,y\in S$ and $\zeta\in\partial(h+\delta_{S})(x)$. By the chain rule theorem, (clarke1990optimization, , Theorem 2.3.10), there exists $\eta\in\partial\varphi(g(x))$ such that $\zeta=J(x)^{T}\eta$. Now, using these properties and the convexity of $\varphi$, we get
$\begin{array}[]{ll}h(y)-h(x)-\langle\zeta,y-x\rangle&=\varphi(g(y))-\varphi(g(x))-\langle J(x)^{T}\eta,y-x\rangle\geq\langle\eta,g(y)-g(x)\rangle-\langle\eta,J(x)(y-x)\rangle\\[5.69054pt] &\geq-\|\eta\|\,\|g(y)-g(x)-J(x)(y-x)\|\geq-\tfrac{L_{0}L_{1}}{\nu+1}\|y-x\|^{\nu+1},\end{array}$
i.e., our desired result follows from Proposition 3.13. ∎
In what follows, we investigate some calculus around the class of $\nu$-paraconvex functions. Since the proofs of these results are straightforward consequence of Definition 3.1, we omit the details of the proofs.
Proposition 3.15 (Paraconvexity calculus).
Let $\mathcal{I}$ be an index set and $S\subseteq\mathbb{R}^{n}$ be a nonempty convex set. Then, the following statements hold:
  1. 1
    Let the function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be $\nu$-paraconvex on $S$ for some $0<\nu\leq 1$ with constant $\rho\geq 0$.
    1. (i)
      If $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a convex function on $S$, then $g+h$ is $\nu$-paraconvex on $S$ with constant $\rho$.
    2. (ii)
      The function $h$ is $\nu$-paraconvex on $S$ with any constant $\varrho\in(\rho,\infty)$.
    3. (iii)
      Given $\varrho>0$, the function $\varrho h$ is $\nu$-paraconvex on $S$ with constant $\varrho\,\rho$.
    4. (iv)
      If $S$ is bounded with diameter $K>0$, then $h$ is $\vartheta-$paraconvex on $S$ with constant $\rho K^{\nu-\vartheta}$ for any $0<\vartheta<\nu$.
  2. 2
    Let $h_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $i\in\mathcal{I}$, be $\nu_{i}-$paraconvex on $S$ for some $0<\nu_{i}\leq 1$ with constant $\rho_{i}\geq 0$.
    1. (i)
      If $\sum_{i}\rho_{i}<\infty$ and $\vartheta=\nu_{i}$, $i\in\mathcal{I},$ then $\sum_{i}h_{i}$ is $\vartheta-$paraconvex on $S$ with constant $\sum_{i}\rho_{i}$.
    2. (ii)
      If $S$ is bounded with diameter $K>0$, $0<\vartheta=\inf_{i}\nu_{i}$, and $\sum_{i}\rho_{i}K^{\nu_{i}-\vartheta}<\infty$, then $\sum_{i}h_{i}$ is $\vartheta-$paraconvex on $S$ with constant $\sum_{i}\rho_{i}K^{\nu_{i}-\vartheta}.$
    3. (iii)
      If $\sup_{i}\rho_{i}<\infty$, and $\vartheta=\nu_{i}$, $i\in\mathcal{I},$ then $\sup_{i}{\mathopen{}\left\{h_{i}(x){}\mid{}{}\right\}\mathclose{}}$ is $\vartheta-$paraconvex on $S$ with constant $\rho:=\sup_{i}\rho_{i}$.
Our subsequent result states that continuous midpoint $\nu$-paraconvex functions are $\nu$-paraconvex, which result simplifies the verification of paraconvexity.
Proposition 3.16 (Midpoint paraconvexity).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuous function on a nonempty convex set $S\subseteq\mathbb{R}^{n}$. Given some $\nu\in(0,1]$, the function $h$ is $\nu$-paraconvex on $S$ if and only if it is midpoint $\nu$-paraconvex on $S$, that is, there exists the constant $\rho\geq 0$ such that
$h\bigg{(}\frac{x+y}{2}\bigg{)}\leq\frac{1}{2}h(x)+\frac{1}{2}h(y)+\frac{\rho}{2}\|x-y\|^{1+\nu},\quad\quad\forall x,y\in S.$ (3.16)
Proof.
The necessity follows directly from Definition 3.1. To justify sufficiency, we assume a constant $\rho\geq 0$ exists such that the inequality 3.16 holds for all $x,y\in S$. Let $x,y\in S$ be arbitrary and constant hereafter. It is clear that the set $\Lambda:=\big{\{}\frac{p}{2^{k}}:k,p\in\mathbb{N},p\leq 2^{k}\big{\}}$ is dense in $[0,1]$. So, it suffices to verify that the inequality 3.1 holds for each $\lambda\in\Lambda$. Particularly, we prove by induction that for each $k\in\mathbb{N}$, the following inequality is satisfied:
$h\bigg{(}\frac{px+qy}{2^{k}}\bigg{)}\leq\frac{p}{2^{k}}h(x)+\frac{q}{2^{k}}h(y)+\rho\bigg{(}\sum_{i=1}^{k}\frac{1}{2^{i}}\bigg{)}\|x-y\|^{1+\nu},$ (3.17)
for all $p,q\in\mathbb{N}$ such that $p+q=2^{k}$. For $k=1$, this inequality holds thanks to 3.16. Suppose 3.17 is satisfied for some $k\in\mathbb{N}$. We show it for $k+1$. Let $p,q\in\mathbb{N}$ such that $p+q=2^{k+1}$. If $p=q=2^{k}$, then there is nothing to prove. Without loss of generality, assume that $p>2^{k}>q$, i.e.,
$h\bigg{(}\frac{px+qy}{2^{k+1}}\bigg{)}$ $=h\Bigg{(}\frac{\frac{(p-2^{k})x+qy}{2^{k}}+x}{2}\Bigg{)}\leq\frac{1}{2}h\bigg{(}\frac{(p-2^{k})x+qy}{2^{k}}\bigg{)}+\frac{1}{2}h(x)+\frac{\rho}{2}{\mathopen{}\left\|\frac{(p-2^{k})x+qy}{2^{k}}-x\right\|\mathclose{}}^{1+\nu}$
$\leq\frac{p}{2^{k+1}}h(x)+\frac{q}{2^{k+1}}h(y)+\rho\bigg{(}\sum_{i=1}^{k+1}\frac{1}{2^{i}}\bigg{)}\|x-y\|^{1+\nu},$
which comes out of the inductive assumption, completing the inductive step. Now, on the basis of the facts that $\Lambda$ is dense in $[0,1]$, $\sum_{i=1}^{\infty}\frac{1}{2^{i}}=1$, and $h$ is a continuous function, it is concluded that
$h(\lambda x+(1-\lambda)y)\leq\lambda h(x)+(1-\lambda)h(y)+\rho\|x-y\|^{1+\nu},$
for each $\lambda\in[0,1]$, completing the proof. ∎
3.2 Paraconvex optimization: Finding global minima
Here, we consider the optimization problem of the form 1.1. Throughout the paper, we consider the following assumption.
Assumption I.
We assume that
  1. 1.
    The set $X\subseteq\mathbb{R}^{n}$ is a nonempty closed and convex set;
  2. 2.
    The set of minimizers $\mathcal{X}^{*}:=\operatorname*{arg\,min}_{x\in X}f(x)$ is nonempty and $f^{*}$ is the optimal value of 1.1;
and the following properties are satisfied for parameters $\rho\geq 0$, $0<\nu\leq 1$, $\tfrac{1}{1+\nu}<\delta\leq 1$ and $\mu>0$:
  1. 3.
    The function $f$ is $\nu$-paraconvex on a open convex set $X^{\prime}\supseteq X$ with constant $\rho$ (see Definition 3.1);
  2. 4.
    The function $f$ admits HEB of order $\delta$ with constant $\mu$.
We note that this class of problems is wide enough to encompass many interesting problems from application domains, as discussed in Section 3.1. In the next example, we verify with a toy example that this class of problems accepts saddle points and global minima can be located in a large basin.
Example 3.17.
Let us consider the function $h:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by $h(x,y)=x^{2}+\Big{(}y^{2}-1\Big{)}^{2}$. We compute
${\nabla}h(x,y)=\begin{bmatrix}2x\\ 4y(y^{2}-1)\end{bmatrix},\quad\quad\nabla^{2}h(x,y)=\begin{bmatrix}2&0\\ 0&-4+12y^{2}\end{bmatrix}.$
It is evident that the function $h$ is not a convex function possessing a saddle point. Specifically, the points $(0,\pm 1)$ and $(0,0)$ are stationary points: $(0,\pm 1)$ are global minima and $(0,0)$ is a saddle point, as indicated by the presence of both positive and negative eigenvalues in the Hessian matrix at $(0,0)$. Furthermore, the function $h$ is a $1$-paraconvex function with constant $\rho>2$. It is also noteworthy that for any global minimum, a ball with radius $r<1$ around the minimum contains no saddle points.
Figure 2: Stationary points of the function $h(x,y)=x^{2}+\Big{(}y^{2}-1\Big{)}^{2}$: global minima at $(0,\pm 1)$ (red points) and a saddle point at $(0,0)$ (white point).
For many problems of the form (1.1), there is a wide basin around the global minima, as suggested in Figure 2. Hence, if a local optimization algorithm initializes from a point in this basin, it should converge fast to a global minimizer. This motivates the quest for developing two-stage algorithms in which (i) the outer algorithm (e.g., a spectral approach) is used to find a reasonable approximation of a global minimizer in the basin of the cost function, and (ii) the inner algorithm is used to achieve fast convergence to a global minimizer. In more detail, some sufficient conditions can be considered under which variants of spectral methods (chi2019nonconvex, , Section VIII) can bring us to the vicinity of a global minimizer, and then we can apply a proper algorithm (e.g., projected subgradient methods; See Section 4) to converge to this global minimizer.
In the next result, we introduce a neighborhood of the optimal solution set of (1.1) that is devoid of extraneous stationary points as a direct result from $\nu$-paraconvexity and Hölderian error boundness. Note that this region ensures a suitable region around global minia such that if a proper algorithm is initialized from there, then it can generate a sequence converting to the global minimizer. In the following result, let us recall that Assumption I holds.
Lemma 3.18.
Let $x\in X$ be a stationary point of 1.1. Then, either $x\in\mathcal{X}^{*}$ or
$\bigg{(}\tfrac{\mu}{\rho}\bigg{)}^{\tfrac{\delta}{\delta(1+\nu)-1}}\leq\operatorname{dist}(x;\mathcal{X}^{*}).$ (3.18)
Proof.
Without loss of generality, let $x\in X\setminus\mathcal{X}^{*}$ be a stationary point of (1.1). Hence, $0\in\partial(f+\delta_{X})(x)$ and Proposition 3.13 yield
$f(y)-f(x)\geq-\rho\|y-x\|^{(1+\nu)},\quad\forall y\in X.$
Setting $\bar{x}\in\operatorname{proj}_{\mathcal{X}^{*}}{(x)}$ and applying HEB, we come to
$\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x;\mathcal{X}^{*})\leq f(x)-f(\bar{x})\leq\rho\|x-\bar{x}\|^{1+\nu}=\rho\operatorname{dist}^{1+\nu}(x;\mathcal{X}^{*}),$
leading to 3.18. ∎
Now, for any $\gamma\in(0,1]$, let us define the tube
$\mathcal{T}_{\gamma}:={\mathopen{}\left\{x\in X:\operatorname{dist}(x;\mathcal{X}^{*})<{\mathopen{}\left(\gamma\tfrac{\mu}{\rho}\right)\mathclose{}}^{\tfrac{\delta}{\delta(1+\nu)-1}}\right\}\mathclose{}},$ (3.19)
which contains no extraneous stationary points of the problem, due to Lemma 3.18. Additionally, we set
$L:=\sup\bigg{\{}\|\zeta\|:\zeta\in\partial f(x),\leavevmode\nobreak\ x\in\mathcal{T}_{1}\bigg{\}}.$
The following lemma provides a key relationship between $\mu$ and $L$. Specifically, for each $x\in\mathcal{T}_{1}$, the ratio $\tau_{x}:=\tau\operatorname{dist}^{\tfrac{1}{\delta}-1}(x;\mathcal{X}^{*})$ lies within the interval $[0,1]$, where $\tau=\tfrac{\mu}{L}$.
Lemma 3.19.
Let $\gamma\in(0,1]$ and $x\in\mathcal{T}_{\gamma}$. Then, $\tau_{x}=\tau\operatorname{dist}^{\tfrac{1}{\delta}-1}(x;\mathcal{X}^{*})\in[0,1]$. In addition, if $\delta=1$, then $\tau_{x}=\tau\in(0,1]$.
Proof.
Let $x\in\mathcal{T}_{\gamma}$. If $x\in\mathcal{X}^{*}$, then trivially $\tau_{x}=0$ and there is nothing to prove, i.e., without loss of generality, we assume that $x\in\mathcal{T}_{\gamma}\setminus\mathcal{X}^{*}$. For $\bar{x}\in\operatorname{proj}_{\mathcal{X}^{*}}{(x)}$, invoking the mean value theorem (mordukhovich2006variational, , Theorem 3.51), there exist $z\in[x,\bar{x})$ and $\zeta\in\partial f(z)$ satisfying $f(x)-f(\bar{x})\leq\langle\zeta,x-\bar{x}\rangle.$ Applying Hölderian property, we come to
$\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x;\mathcal{X}^{*})\leq f(x)-f(\bar{x})\leq\|\zeta\|\leavevmode\nobreak\ \|x-\bar{x}\|=\|\zeta\|\leavevmode\nobreak\ \operatorname{dist}(x;\mathcal{X}^{*}).$
Furthermore, it is easy to see that $z\in\mathcal{T}_{\gamma}\subseteq\mathcal{T}_{1}$, and so $\|\zeta\|\leq L$. Thus, $\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x;\mathcal{X}^{*})\leq L\operatorname{dist}(x;\mathcal{X}^{*})$, and consequently $\tau_{x}\leq 1$. In the special case where $\delta=1$, it follows that $\tau_{x}=\tau=\tfrac{\mu}{L}\neq 0$. ∎
4 Projected subgradient algorithm for paraconvex optimization
In this section, we present the projected subgradient methods for the nonsmooth and constrained paraconvex optimization problems of the form 1.1 and establish their convergence analysis for several choices of step-sizes. Let us begin with the following generic projected subgradient method where the starting point $x_{0}$ satisfies
$x_{0}\in\mathcal{T}_{\gamma},$
as $\mathcal{T}_{\gamma}$ is defined in (3.19).
Input: $x_{0}\in\mathcal{T}_{\gamma}$$\gamma\in(0,1]$$\alpha_{0}>0$.
1 begin
2       while the stopping criteria does not hold do
3             Choose $\zeta_{k}\in\partial f(x_{k})$;
4
5            Set $x_{k+1}=\operatorname{proj}_{X}{\Big{(}x_{k}-\alpha_{k}\tfrac{\zeta_{k}}{\|\zeta_{k}\|}\Big{)}}$ and $k=k+1$;
6
7      Set $x_{k+1}=x_{k}$;
8
9
Output: $x_{k+1}$.
Algorithm 1 Generic Projected Subgradient Method
In the above algorithm, the step-size $\alpha_{k}>0$ plays a key role in the algorithm’s progression. In fact, by employing different step-sizes, such as fixed, diminishing, and Scaled Polyak’s step-sizes, we can define various projected subgradient methods, where they can then be compared in terms of the convergence rates and overall numerical performance.
We begin with the subsequent lemma establishing a fundamental recurrence, providing an upper bound for the update step based on the chosen step-size. This result is instrumental in deriving the convergence rate of the projected subgradient methods.
Lemma 4.1 (Basic inequalities I).
Let $x_{k}$ and $x_{k+1}$ be two consecutive iterations of Algorithm 1 such that $x_{k}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$. Then, one has
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\tfrac{\alpha_{k}}{L}(f(x_{k})-f^{*})+\alpha_{k}^{2},\hskip 8.53581pt$ (4.1)
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\alpha_{k}\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})+\alpha_{k}^{2}.$ (4.2)
If, in addition, $\alpha_{k}\leq\min\Big{\{}1,\min\{1,\tau\}(\tfrac{\mu}{2\rho})^{\tfrac{\delta}{\delta(1+\nu)-1}}\Big{\}}$, then $x_{k+1}\in\mathcal{T}_{\gamma}$.
Proof.
For Algorithm 1, if $\zeta_{k}=0$, the claims are evident then from $x_{k+1}=x_{k}\in\mathcal{X}^{*}$. Without loss of generality, we assume $\zeta_{k}\neq 0$. To show 4.1, let $x^{*}\in\operatorname{proj}_{\mathcal{X}^{*}}(x_{k})$. Then, $\operatorname{dist}(x_{k};\mathcal{X}^{*})=\|x_{k}-x^{*}\|$ and $f(x^{*})=f^{*}$. Applying on the nonexpansiveness of $\operatorname{proj}_{X}$, it holds that
$\begin{array}[]{ll}\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})&\leq\|x_{k+1}-x^{*}\|^{2}\leq\|(x_{k}-x^{*})-\alpha_{k}\tfrac{\zeta_{k}}{\|\zeta_{k}\|}\|^{2}=\|x_{k}-x^{*}\|^{2}+\tfrac{2\alpha_{k}}{\|\zeta_{k}\|}\langle\zeta_{k},x^{*}-x_{k}\rangle+\alpha_{k}^{2}\\[5.69054pt] &\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{2\alpha_{k}}{\|\zeta_{k}\|}{\mathopen{}\left(f^{*}-f(x_{k})+\rho\operatorname{dist}^{1+\nu}(x_{k};\mathcal{X}^{*})\right)\mathclose{}}+\alpha_{k}^{2},\end{array}$ (4.3)
where the last inequality comes from $\nu$-paraconvexity of $f$, Proposition 3.13 (i). Now, inasmuch as $x_{k}\in\mathcal{T}_{\gamma}\subseteq\mathcal{T}_{1}$, we have $\tfrac{\alpha_{k}}{L}\leq\tfrac{\alpha_{k}}{\|\zeta_{k}\|}$ and $\operatorname{dist}(x_{k};\mathcal{X}^{*})<{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{\delta}{\delta(1+\nu)-1}}$, i.e.,
$\rho\operatorname{dist}^{1+\nu}(x_{k};\mathcal{X}^{*})<{\tfrac{1}{2}}\,\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})\leq{\tfrac{1}{2}}(f(x_{k})-f^{*}),$
using Hölderian property of $f$. Substituting these bounds into 4.3 results in 4.1, i.e.,
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{2\alpha_{k}}{\|\zeta_{k}\|}\Big{(}f^{*}-f(x_{k})+{\tfrac{1}{2}}(f(x_{k})-f^{*})\Big{)}+\alpha_{k}^{2}\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\tfrac{\alpha_{k}}{L}{\mathopen{}\left(f(x_{k})-f^{*}\right)\mathclose{}}+\alpha_{k}^{2}.$
Moreover, the argument for 4.2 is derived from 4.1, relying on Hölderian property of $f$, i.e.,
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\tfrac{\alpha_{k}}{L}{\mathopen{}\left(f(x_{k})-f^{*}\right)\mathclose{}}+\alpha_{k}^{2}\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\alpha_{k}\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})+\alpha_{k}^{2}.$
For deriving the last part, $x_{k+1}\in\mathcal{T}_{\gamma}$, we consider two cases: (i) $\alpha_{k}\leq\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})$; (ii) $\alpha_{k}>\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})$.
In Case (i), $\alpha_{k}\leq\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})$, applying the above inequality to 4.1 yields
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\alpha_{k}\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})+\alpha_{k}^{2}\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})<{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{1-\delta(1+\nu)}}.$
i.e., $x_{k+1}\in\mathcal{T}_{\gamma}$.
For Case (ii), $\alpha_{k}>\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})$, leading to
$\operatorname{dist}(x_{k};\mathcal{X}^{*})<{\mathopen{}\left(\tfrac{\alpha_{k}}{\tau}\right)\mathclose{}}^{\delta}.$ (4.4)
Take the convex function $\varphi(t):=t-\alpha_{k}\tau t^{\tfrac{1}{2\delta}}+\alpha_{k}^{2}$ on $[0,(\tfrac{\alpha_{k}}{\tau})^{2\delta}]$ where the maximum is achieved at 00 or $(\tfrac{\alpha_{k}}{\tau})^{2\delta}$. On the basis of 4.1 and 4.4, $\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\varphi(\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*}))$ and $\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})\in[0,(\tfrac{\alpha_{k}}{\tau})^{2\delta}]$, i.e.,
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\max{\mathopen{}\left\{\varphi(t):\leavevmode\nobreak\ t\in[0,(\tfrac{\alpha_{k}}{\tau})^{2\delta}]\right\}\mathclose{}}=\max{\mathopen{}\left\{\alpha^{2}_{k},\leavevmode\nobreak\ (\tfrac{\alpha_{k}}{\tau})^{2\delta}-\alpha_{k}\tau\tfrac{\alpha_{k}}{\tau}+\alpha^{2}_{k}\right\}\mathclose{}}\leq\max{\mathopen{}\left\{\alpha^{2\delta}_{k},\leavevmode\nobreak\ (\tfrac{\alpha_{k}}{\tau})^{2\delta}\right\}\mathclose{}}$
$\leq\max{\mathopen{}\left\{1,\leavevmode\nobreak\ \tfrac{1}{\tau^{2\delta}}\right\}\mathclose{}}\min\{1,\tau^{2\delta}\}(\tfrac{\mu}{2\rho})^{\tfrac{2\delta}{\delta(1+\nu)-1}}=(\tfrac{\mu}{2\rho})^{\tfrac{2\delta}{\delta(1+\nu)-1}},$
where the second and third inequalities are derived from the bound on $\alpha_{k}$, i.e., $x_{k+1}\in\mathcal{T}_{\gamma}$. ∎
Next, we drive an upper bound for the difference between the lowest value of the function over the iterations and the optimal value.
Lemma 4.2 (Basic inequalities II).
Given the iterations $\{x_{k}\}_{k\geq 0}\subseteq\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$, generated by Algorithm 1, one has
$f(x_{k})-f^{*}\leq\frac{L\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+L\alpha_{k}^{2}}{\alpha_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \,$ (4.5)
$f_{k}^{*}-f^{*}\leq\frac{L\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})+L\sum_{i=1}^{k}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}},$ (4.6)
where $f_{k}^{*}:=\min\{f(x_{i}):i=1,2,\ldots,k\}$.
Proof.
The inequality 4.5 follows directly from Lemma 4.1, inasmuch as $\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\geq 0$. Let us justify 4.6. From the inequality 4.1 in Lemma 4.1 and $f^{*}_{k}\leq f(x_{i})$, we obtain
$\operatorname{dist}^{2}(x_{i+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{i};\mathcal{X}^{*})-\tfrac{\alpha_{i}}{L}(f^{*}_{k}-f^{*})+\alpha_{i}^{2},$
for all $i=1,2,\ldots,k$. Summing this inequality from $i=1$ to $k$, it holds that
$\sum_{i=1}^{k}\operatorname{dist}^{2}(x_{i+1};\mathcal{X}^{*})\leq\sum_{i=1}^{k}\operatorname{dist}^{2}(x_{i};\mathcal{X}^{*})-\sum_{i=1}^{k}\tfrac{\alpha_{i}}{L}(f^{*}_{k}-f^{*})+\sum_{i=1}^{k}\alpha_{i}^{2},$
leading to
$0\leq\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})-\tfrac{f^{*}_{k}-f^{*}}{L}\sum_{i=1}^{k}\alpha_{i}+\sum_{i=1}^{k}\alpha_{i}^{2},$
resulting in 4.6. ∎
4.1 Projected subgradient method with constant step-size
Here, we analyze a projected subgradient method with a constant step-size, $\alpha_{k}=\alpha>0$ for all $k$. We show that, with appropriate initialization, the sequence $D_{k}:=\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})$ converges linearly below a certain fixed threshold.
Theorem 4.3 (Convergence of constant step-size method).
Let $\tau=\tfrac{\mu}{L}$. Fix a constant step-size $\alpha$ satisfying
[…]
Let $\{x_{k}\}_{k\geq 0}$ be the sequence of iterations generated by Algorithm 1 with the constant step-size $\alpha_{k}=\alpha$, let the initial point $x_{0}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$, and let
$D_{*}:={\mathopen{}\left(\tfrac{\alpha}{\tau}\right)\mathclose{}}^{2\delta},\quad\quad q:=1-\tfrac{\alpha\tau}{2\delta}{\mathopen{}\left(\tfrac{2\rho}{\mu}\right)\mathclose{}}^{\tfrac{2\delta-1}{\delta(1+\nu)-1}},\quad\quad\mathcal{D}^{2}:=\max{\mathopen{}\left\{D_{0},\alpha^{2}+D_{*}\right\}\mathclose{}}.$
Then, $10<q<$ 1, and for any $k\in\mathbb{N}$, we have
$\sqrt{D_{k}}\leq\mathcal{D}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{\delta}{\delta(1+\nu)-1}},\quad\text{ and }\quad D_{k}-D_{*}\leq\max{\mathopen{}\left\{q^{k}{\mathopen{}\left(D_{0}-D_{*}\right)\mathclose{}},\alpha^{2}\right\}\mathclose{}}.$ (4.7)
Proof.
We first verify claims about $\mathcal{D}$ and $q$, and then use the induction to establish the iteration-dependent inequalities. First, we will verify $\mathcal{D}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{\delta}{\delta(1+\nu)-1}}$. Inasmuch as $x_{0}\in\mathcal{T}_{\gamma}$, $D_{0}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}}$ due to definition of $\mathcal{T}_{\gamma}$. Moreover, using the upper bound of $\alpha$, we get
$\alpha^{2}+D_{*}\leq\alpha^{2\delta}+\tfrac{\alpha^{2\delta}}{\tau^{2\delta}}=\alpha^{2\delta}\tfrac{\tau^{2\delta}+1}{\tau^{2\delta}}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(\nu+1)-1}},$
i.e., $\mathcal{D}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}}.$ Regarding to $q\in(0,1)$, it is clear that $q<1$. Considering
$0<\alpha<\tfrac{2\delta}{\tau}{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta-1}{\delta(\nu+1)-1}},$
and applying the basic calculus, it holds that $q>0$. We now prove the inequalities in 4.7 using induction. At $k=0$, the inequalities hold trivially. Now by induction, assume that for some $k\in\mathbb{N}$, the following holds (inductive hypothesis):
$\sqrt{D_{i}}\leq\mathcal{D}\,\,\text{ and }\,\,D_{i}-D_{*}\leq\max{\mathopen{}\left\{q^{i}{\mathopen{}\left(D_{0}-D_{*}\right)\mathclose{}},\alpha^{2}\right\}\mathclose{}},\quad\forall i=0,1,\ldots,k.$
We will verify these claims for $(k+1)-$th step. From Lemma 4.1, and the concavity of function $t^{\tfrac{1}{2\delta}}$ on $[0,+\infty)$, it follows that
$D_{k+1}\leq D_{k}-\alpha\tau D_{k}^{\tfrac{1}{2\delta}}+\alpha^{2}=D_{k}-\alpha\tau{\mathopen{}\left(D_{k}^{\tfrac{1}{2\delta}}-\tfrac{\alpha}{\tau}\right)\mathclose{}}=D_{k}-\alpha\tau{\mathopen{}\left(D_{k}^{\tfrac{1}{2\delta}}-D_{*}^{\tfrac{1}{2\delta}}\right)\mathclose{}}\leq D_{k}-\tfrac{\alpha\tau}{2\delta D_{k}^{1-\tfrac{1}{2\delta}}}(D_{k}-D_{*}),$
i.e.,
$D_{k+1}-D_{*}\leq\bigg{(}1-\tfrac{\alpha\tau}{2\delta D_{k}^{1-\tfrac{1}{2\delta}}}\bigg{)}(D_{k}-D_{*}).$ (4.8)
We have two cases: (i) $D_{k}\geq D_{*}$; (ii) $D_{k}<D_{*}$.
In Case (i), since $D_{k}\leq\mathcal{D}^{2}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}}$, we obtain
$D_{k+1}-D_{*}\leq\bigg{(}1-\tfrac{\alpha\tau}{2\delta D_{k}^{1-\tfrac{1}{2\delta}}}\bigg{)}(D_{k}-D_{*})\leq\bigg{(}1-\tfrac{\alpha\tau}{2\delta{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}\tfrac{2\delta-1}{2\delta}}}\bigg{)}(D_{k}-D_{*})=q{\mathopen{}\left(D_{k}-D_{*}\right)\mathclose{}}.$
Moreover, $\sqrt{D_{k+1}}\leq\sqrt{D_{k}}\leq\mathcal{D}$, inasmuch as $10<q<$ 1. In Case (ii), $D_{k}<D_{*}$, relying once again on Lemma 4.1, we get
$D_{k+1}-D_{*}\leq D_{k}-\alpha\tau D_{k}^{\tfrac{1}{2\delta}}+\alpha^{2}-D_{*}\leq D_{k}+\alpha^{2}-D_{*}<\alpha^{2}.$
In addition, $\sqrt{D_{k+1}}\leq\sqrt{\alpha^{2}+D_{*}}\leq\mathcal{D}$. Hence, in both cases and by the induction assumption, it can be deduced
$D_{k+1}-D_{*}\leq\max{\mathopen{}\left\{q{\mathopen{}\left(D_{k}-D_{*}\right)\mathclose{}},\alpha^{2}+D_{*}\right\}\mathclose{}}\leq\max{\mathopen{}\left\{q^{k+1}{\mathopen{}\left(D_{0}-D_{*}\right)\mathclose{}},\alpha^{2}+D_{*}\right\}\mathclose{}}.$
Since $\sqrt{D_{k+1}}\leq\mathcal{D}$, by induction, the claims hold for all $k\geq 0$. ∎
Theorem 4.3 establishes that the sequence $D_{k}$ decrease to a value below $D_{*}$ at a linear rate. Furthermore, even if $D_{k}$ becomes smaller than $D_{*}$, it remains close to this threshold.
The subsequent result is a straightforward consequence of Theorem 4.3 for $\delta=1$ under Assumption I.
Corollary 4.4.
Let $\delta=1$ and $\tau=\tfrac{\mu}{L}$. Fix a constant step-size $\alpha$ satisfying
[…]
Let $\{x_{k}\}_{k\geq 0}$ be the sequence of iterations generated by Algorithm 1 with the constant step-size $\alpha_{k}=\alpha$ and let the initial point $x_{0}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$. For
$D_{*}:=\tfrac{\alpha^{2}}{\tau^{2}},\quad\quad q:=1-\tfrac{\alpha\tau}{2}{\mathopen{}\left(\tfrac{2\rho}{\mu}\right)\mathclose{}}^{\tfrac{1}{\nu}},\quad\quad\mathcal{D}^{2}:=\max{\mathopen{}\left\{D_{0},\alpha^{2}+D_{*}\right\}\mathclose{}},0<q<1,$
and for any $k\in\mathbb{N}$, then
$\sqrt{D_{k}}\leq\mathcal{D}\leq{\mathopen{}\left(\dfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{1}{\nu}},\quad\text{ and }\quad D_{k}-D_{*}\leq\max{\mathopen{}\left\{q^{k}{\mathopen{}\left(D_{0}-D_{*}\right)\mathclose{}},\alpha^{2}\right\}\mathclose{}}.$
Proof.
Using Lemma 4.1, it is clear that $\tau\in(0,1)$. By straightforward calculations, the result follows directly from Theorem 4.3. ∎
The next theorem provides an upper bound, determined by the step-size, for the distance of the lowest estimated value over iterations from the optimal value.
Theorem 4.5 (Convergence rate of constant step-size method).
Let $\tau=\tfrac{\mu}{L}$. Fix a constant step-size $\alpha$ satisfying
$0<\alpha<\min{\mathopen{}\left\{1,\,\min\{1,\tau\}(\tfrac{\mu}{2\rho})^{\tfrac{1}{\delta(1+\nu)-1}}\right\}\mathclose{}}.$
Let $\{x_{k}\}_{k\geq 1}$ be the sequence of iterations generated by Algorithm 1 with the constant step-size $\alpha_{k}=\alpha$ and let the initial point $x_{1}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$. Then,
$0\leq f^{*}_{k}-f^{*}\leq 2L\alpha,\quad\quad\forall k\geq\tfrac{1}{\alpha^{2}}{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}},$
where $f_{k}^{*}:=\min\{f(x_{i}):i=1,2,\ldots,k\}$.
Proof.
By Lemma 4.1, the upper bounds of $\alpha$ ensures that $x_{k}\in\mathcal{T}_{\gamma}$, for each $k$, i.e.,
$f_{k}^{*}-f^{*}\leq\frac{L\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})+L\sum_{i=1}^{k}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}\leq\frac{L{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}}+Lk\alpha^{2}}{k\alpha}\leq\frac{2Lk\alpha^{2}}{k\alpha}=2L\alpha,$
where the first inequality due to Lemma 4.2. ∎
4.2 Projected subgradient methods with diminishing step-sizes
In the previous section, we analyzed the constant step-size scheme and established convergence results within a fixed threshold. Here, we shift our focus to achieving convergence to the optimal solution by studying the projected subgradient method with a diminishing step-size. Specifically, we examine the convergence behavior of the method under two distinct step-size strategies: a nonsummable diminishing step-size and a square-summable yet nonsummable step-size.
The next result validates the convergence of the sequence $f^{*}_{k}=\min\{f(x_{i}):i=1,2,\ldots,k\}$ in which the iterations $x_{k}$ are generated by the projected subgradient method with nonsummable diminishing step-size satisfying
$\alpha_{k}\geq 0,\quad\lim_{k\to\infty}\alpha_{k}=0,\quad\sum_{k=1}^{\infty}\alpha_{k}=\infty.$
Furthermore, it verifies the convergence of a subsequence of the iterations $x_{k}$.
Theorem 4.6 (Convergence of nonsummable diminishing method).
Let $\tau=\tfrac{\mu}{L}$ and let $\alpha_{k}$ be nonsummable diminishing step-size satisfying
$0\leq\alpha_{k}<\min{\mathopen{}\left\{1,\,\min\{1,\tau\}(\tfrac{\mu}{2\rho})^{\tfrac{1}{\delta(1+\nu)-1}}\right\}\mathclose{}}.$
Let $\{x_{k}\}_{k\geq 1}$ be the iterations generated by Algorithm 1 with step-size $\alpha_{k}$ and let the initial point $x_{1}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$. Then,
$\lim_{k\to\infty}f^{*}_{k}=f^{*},\quad\text{and}\quad\liminf_{k\to\infty}f(x_{k})=f^{*}.$
Moreover, if $\{x_{k}\}_{k\geq 1}$ is a bounded sequence, it has a convergent subsequence to some optimal solution $x^{*}\in\mathcal{X}^{*}$.
Proof.
According to Lemma 4.1, we have $x_{k}\in\mathcal{T}_{\gamma}$ for all $k\geq 1$, i.e.,
$0\leq f_{k}^{*}-f^{*}\leq\frac{L\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})+L\sum_{i=1}^{k}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}},$
due to Lemma 4.2. Let us consider a fixed $j\in\mathbb{N}$. For $k>j$, it holds that
$0\leq\frac{\sum_{i=1}^{k}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}=\frac{\sum_{i=1}^{j-1}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}+\frac{\sum_{i=j}^{k}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}\leq\frac{\sum_{i=1}^{j-1}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}+\displaystyle\max_{j\leq i\leq k}\{\alpha_{i}\}\frac{\sum_{i=j}^{k}\alpha_{i}}{\sum_{i=1}^{k}\alpha_{i}}\leq\frac{\sum_{i=1}^{j-1}\alpha_{i}^{2}}{\sum_{i=1}^{k}\alpha_{i}}+\displaystyle\max_{j\leq i}\{\alpha_{i}\}.$
Taking the limit as $k\to\infty$, we get $0\leq\tfrac{\sum_{i=1}^{\infty}\alpha_{i}^{2}}{\sum_{i=1}^{\infty}\alpha_{i}}\leq\max_{j\leq i}\{\alpha_{i}\}$, for each $j\in\mathbb{N}$. Now, taking limit as $j\to\infty$, we conclude that $\tfrac{\sum_{i=1}^{\infty}\alpha_{i}^{2}}{\sum_{i=1}^{\infty}\alpha_{i}}=0$. On the other hand, $\tfrac{L\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})}{\sum_{i=1}^{\infty}\alpha_{i}}=0$. Thus, $\lim_{k\to\infty}f_{k}^{*}=f^{*}$.
We next demonstrate that $\liminf_{k\to\infty}f(x_{k})=f^{*}.$ To this end, we construct a subsequence $\{x_{k_{j}}\}_{j\in\mathbb{N}}$ of $\{x_{k}\}_{k\in\mathbb{N}}$ converging to some optimal solution. If $f(x_{k})=f^{*}$ for some $k\in\mathbb{N}$, the result is valid. Otherwise, assume $f^{*}<f(x_{k})$, for all $k\in\mathbb{N}$. For each $k\in\mathbb{N}$, there exists nondecreasing index $i_{k}\in\{1,2,\ldots,k\}$ such that $f^{*}_{k}=f(x_{i_{k}})$. Therefore, we have $\lim_{k\to\infty}f(x_{i_{k}})=f^{*}$. Let us define $k_{1}:=i_{1}=1$. Since, $f^{*}<f(x_{1})$, there exists $k\in\mathbb{N}$, such that $f(x_{i_{k}})<f(x_{1})$. Now, let us set $k_{2}:=\min\{i_{k}\in\mathbb{N}:f(x_{i_{k}})<f(x_{1})\}$, i.e., $k_{1}=1<k_{2}$ and $f(x_{k_{2}})<f(x_{k_{1}})$. Continuing this process, we construct a subsequence $x_{k_{j}}$ such that
$k_{1}<k_{2}<\ldots<k_{j},\quad\text{and}\quad f(x_{k_{j}})<f(x_{k_{j-1}})<\ldots<f(x_{k_{1}}),$
where $k_{t}=\min\{i_{k}\in\mathbb{N}:f(x_{i_{k}})<f(x_{k_{t-1}})\}$ for each $1\leq t\leq j$. Now, noticing $\lim_{k\to\infty}f(x_{i_{k}})=f^{*}<f(x_{k_{j}})$, there exists $k\in\mathbb{N}$ such that $f(x_{i_{k}})<f(x_{k_{j}})$. As such, setting $k_{j+1}:=\min\{i_{k}\in\mathbb{N}:f(x_{i_{k}})<f(x_{k_{j}})\}$, we have $f(x_{k_{j+1}})<f(x_{k_{j}})$ and $k_{j}<k_{j+1}$. Otherwise, if $k_{j+1}<k_{j}$, $f(x_{k_{j+1}})<f(x_{k_{j}})<f(x_{k_{j-1}})$, this leads to a contradiction. Thus, $\{x_{k_{j}}\}_{j\in\mathbb{N}}$ is a subsequence of sequences $\{x_{k}\}_{k\in\mathbb{N}}$ and $\{x_{i_{k}}\}_{k\in\mathbb{N}}$. Therefore, $\lim_{j\to\infty}f(x_{k_{j}})=\displaystyle\lim_{k\to\infty}f(x_{i_{k}})=f^{*}$ and
$f^{*}\leq\liminf_{k\to\infty}f(x_{k})\leq\lim_{j\to\infty}f(x_{k_{j}})=f^{*}.$
This implies $\liminf_{k\to\infty}f(x_{k})=f^{*}$.
We address the final claim by assuming that $\{x_{k}\}_{k\in\mathbb{N}}$ is a bounded sequence. Consequently, $\{x_{k_{j}}\}_{j\in\mathbb{N}}$ is also bounded, implying that it has a convergent subsequence. Without loss of generality, suppose $x_{k_{j}}\to x^{*}$ for some $x^{*}\in X$, i.e.,
$f^{*}=\lim_{j\to\infty}f(x_{k_{j}})=f(x^{*}).$
As such, $x^{*}\in\mathcal{X}^{*}$, completing the proof. ∎
The previous theorem established that projected subgradient methods with nonsummable step-size exhibit subsequential convergence. The following theorem extends this result by demonstrating the convergence of the iterations $x_{k}$ generated by the projected subgradient method when the step-size satisfies the following conditions:
$\alpha_{k}\geq 0,\quad\sum_{k=1}^{\infty}\alpha_{k}=\infty,\quad\sum_{k=1}^{\infty}\alpha_{k}^{2}<\infty.$
Theorem 4.7 (Convergence of square-summable but not summable method).
Let $\tau=\tfrac{\mu}{L}$ and let $\alpha_{k}$ be the square-summable but not summable step-size satisfying
$0\leq\alpha_{k}<\min{\mathopen{}\left\{1,\,\min\{1,\tau\}(\tfrac{\mu}{2\rho})^{\tfrac{1}{\delta(1+\nu)-1}}\right\}\mathclose{}}.$
Let $\{x_{k}\}_{k\geq 1}$ be the iterations generated by Algorithm 1 with step-size $\alpha_{k}$ and let the initial point $x_{1}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$. Then,
$\lim_{k\to\infty}f(x_{k})=\lim_{k\to\infty}f^{*}_{k}=f^{*},$
and the sequence of iterations $\{x_{k}\}_{k\geq 1}$ converges to some optimal solution $x^{*}\in\mathcal{X}^{*}$.
Proof.
Invoking Theorem 4.6, it is clear that $\liminf_{k\to\infty}f(x_{k})=\lim_{k\to\infty}f^{*}_{k}=f^{*},$ which ensures that
$\lim_{j\to\infty}f(x_{k_{j}})=f^{*},$ (4.9)
for some subsequence $\{x_{k_{j}}\}_{j\in\mathbb{N}}$ of $\{x_{k}\}_{k\in\mathbb{N}}$. To establish the convergence of $x_{k}$ to some optimal solution, we show that $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ tends to zero. By Lemma 4.1, for each $k\in\mathbb{N}$, we have $x_{k}\in\mathcal{T}_{\gamma}$, and
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\tfrac{\alpha_{k}}{L}{\mathopen{}\left(f(x_{k})-f^{*}\right)\mathclose{}}+\alpha_{k}^{2}\leq\operatorname{dist}^{2}(x_{k-1};\mathcal{X}^{*})+\alpha_{k-1}^{2}+\alpha_{k}^{2}$
$\leq\ldots\leq\operatorname{dist}^{2}(x_{1};\mathcal{X}^{*})+\sum_{i=1}^{k}\alpha_{i}^{2}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{2\delta}{\delta(1+\nu)-1}}+\sum_{i=1}^{\infty}\alpha_{i}^{2}.$ (4.10)
As such, $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ is a bounded sequence, inasmuch as $\sum_{i=1}^{\infty}\alpha_{i}^{2}<\infty$, i.e., $\operatorname{dist}(x_{k_{j}};\mathcal{X}^{*})$ has a convergent subsequence. Without loss of generality, we assume that $\operatorname{dist}(x_{k_{j}};\mathcal{X}^{*})$ converges to some $d\geq 0$, and hence we need to show $d=0$. Suppose on the contrary that $d>0$, i.e., for sufficiently large $j$, $\operatorname{dist}(x_{k_{j}};\mathcal{X}^{*})>\tfrac{d}{2}$. Using this inequality and the Hölderian property, it follows that
$\mu{\mathopen{}\left(\tfrac{d}{2}\right)\mathclose{}}^{\frac{1}{\delta}}<\mu\operatorname{dist}^{\frac{1}{\delta}}(x_{k_{j}};\mathcal{X}^{*})\leq f(x_{k_{j}})-f^{*},$
for sufficiently large $j$, which contradicts 4.9, i.e., $\operatorname{dist}(x_{k_{j}};\mathcal{X}^{*})\to 0$ as $j\to\infty$. Let $\epsilon>0$ be arbitrary and fixed. There exists $j_{0}\in\mathbb{N}$ such that $\operatorname{dist}(x_{k_{j_{0}}};\mathcal{X}^{*})<\tfrac{\epsilon^{2}}{2}$ and $\sum_{i=k_{j_{0}}}^{k}\alpha_{i}^{2}<\tfrac{\epsilon^{2}}{2}$, for each $k\geq k_{j_{0}}$. Following 4.10, for each $k\geq k_{j_{0}}$, yields
$\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k_{j_{0}}};\mathcal{X}^{*})+\sum_{i=k_{j_{0}}}^{k}\alpha_{i}^{2}<\tfrac{\epsilon^{2}}{2}+\tfrac{\epsilon^{2}}{2}=\epsilon^{2}.$
Consequently, $\operatorname{dist}(x_{k};\mathcal{X}^{*})\to 0$. On the other hand, since $\mathcal{X}^{*}$ is closed, $x_{k}\to x^{*}$ for some $x^{*}\in\mathcal{X}^{*}$. In addition, we get $\lim_{k\to\infty}f(x_{k})=f(x^{*})=f^{*},$ adjusting our desired results. ∎
We next analyze the convergence of the projected subgradient methods with specific step-size schemes, including diminishing step-sizes and geometrically decaying step-size. First, we focus on the projected subgradient method with diminishing step-size of the form, $\alpha_{k}=\lambda(k+\beta)^{-r}$ where $\lambda,\beta>0$ and $10<r<$ 1. Such step-sizes are widely used in both stochastic and deterministic versions of the subgradient method. The sublinear convergence of the projected subgradient method with this step-size is established in the following result.
Theorem 4.8 (Convergence rate of diminishing method).
Let $r\in(0,1)$ and $\lambda>0$. Set
$A:=2^{r\delta}\leavevmode\nobreak\ \sqrt{{\mathopen{}\left(\tfrac{2\lambda}{\tau}\right)\mathclose{}}^{2\delta}+\lambda^{2}},\quad\text{and}\quad\beta:=\max{\mathopen{}\left\{{\mathopen{}\left(\tfrac{4r\delta A^{\tfrac{2\delta-1}{\delta}}}{\lambda\tau}\right)\mathclose{}}^{\tfrac{1}{1-2r(1-\delta)}},A^{\tfrac{1}{\delta r}}{\mathopen{}\left(\tfrac{2\rho}{\mu}\right)\mathclose{}}^{\tfrac{1}{r(\delta(1+\nu)-1)}}\right\}\mathclose{}}.$
Let the iterations $\{x_{k}\}_{k\geq 0}$ be generated by Algorithm 1 with step-size $\alpha_{k}=\lambda(k+\beta)^{-r}$ and let the initial point $x_{0}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$ satisfying $\operatorname{dist}(x_{0};\mathcal{X}^{*})\leq\min{\mathopen{}\left\{(\tfrac{2\lambda}{\tau})^{\delta},A\beta^{-\delta r}\right\}\mathclose{}}$. Then,
$\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq\dfrac{A}{(k+\beta)^{\delta r}}.$ (4.11)
Proof.
Clearly, $\operatorname{dist}(x_{0};\mathcal{X}^{*})\leq A\beta^{-\delta r}$, by virtue of the initialization condition. Let us proceed by induction. Assume that the inequality 4.11 holds for $k$, $\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq A(k+\beta)^{-\delta r}$. Noticing $\beta\geq A^{\tfrac{1}{\delta r}}{\mathopen{}\left(\tfrac{2\rho}{\mu}\right)\mathclose{}}^{\tfrac{1}{r(\delta(1+\nu)-1)}}$, it follows that $x_{k}\in\mathcal{T}_{\gamma}$. From Lemma 4.1, we obtain
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\lambda(k+\beta)^{-r}\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})+\lambda^{2}(k+\beta)^{-2r}.$ (4.12)
Let us consider the index set
$I:={\mathopen{}\left\{i\in\mathbb{N}:\leavevmode\nobreak\ {\tfrac{1}{2}}\,\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{i};\mathcal{X}^{*})\leq\frac{\lambda}{(i+\beta)^{r}}\right\}\mathclose{}}.$
There are two possible cases for the index $k+1$: (i) $k+1\in I$; (ii) $k+1\not\in I$. In Case (i), $k+1\in I$, it follows that: $\operatorname{dist}(x_{k+1};\mathcal{X}^{*})\leq\frac{A}{(k+1+\beta)^{\delta r}}$, since ${\mathopen{}\left(\tfrac{2\lambda}{\tau}\right)\mathclose{}}^{\delta}<A$. Alternatively, in Case (ii), $k+1\not\in I$, let us further consider two subcases: $k$: either $k\in I$ or $k\not\in I$. In the first case, $k\in I$, employing the inequality 4.12, we come to
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\frac{\lambda^{2}}{(k+\beta)^{2r}}\leq\frac{{\mathopen{}\left(\tfrac{2\lambda}{\tau}\right)\mathclose{}}^{2\delta}}{(k+\beta)^{2\delta r}}+\frac{\lambda^{2}}{(k+\beta)^{2r}}\leq{\mathopen{}\left({\mathopen{}\left(\tfrac{2\lambda}{\tau}\right)\mathclose{}}^{2\delta}+\lambda^{2}\right)\mathclose{}}\frac{1}{(k+\beta)^{2\delta r}}\leq\frac{A^{2}}{(k+1+\beta)^{2\delta r}}.$
We now verify the second case, $k\not\in I$, i.e.,
${\tfrac{1}{2}}\,\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})>\dfrac{\lambda}{(k+\beta)^{r}},$
Substituting this into 4.12 yields
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})<\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\dfrac{{\tfrac{1}{2}}\lambda\tau}{(k+\beta)^{r}}\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*}).$ (4.13)
Let us consider the convex function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ given by
$\varphi(t):=t-\dfrac{{\tfrac{1}{2}}\lambda\tau}{(k+\beta)^{r}}t^{\tfrac{1}{2\delta}},$
defined over $[0,A^{2}(k+\beta)^{-2\delta r}]$ where the function achieves its maximum at 00 or $A^{2}(k+\beta)^{-2\delta r}$. Using the inductive assumption and (4.12), we deduce
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})<\varphi(\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})\leq\max{\mathopen{}\left\{\varphi(0),\varphi(A^{2}(k+\beta)^{-2\delta r})\right\}\mathclose{}}=\dfrac{A^{2}}{(k+\beta)^{2\delta r}}-\dfrac{{\tfrac{1}{2}}\lambda\tau A^{\tfrac{1}{\delta}}}{(k+\beta)^{2r}}.$
To ensure the result, it is enough to verify
$\dfrac{A^{2}}{(k+\beta)^{2\delta r}}-\dfrac{\tfrac{1}{2}\lambda\tau A^{\tfrac{1}{\delta}}}{(k+\beta)^{2r}}\leq\dfrac{A^{2}}{(k+1+\beta)^{2\delta r}}.$ (4.14)
Using the convexity of the function $\dfrac{A^{2}}{x^{2\delta r}}$ on positive real number set, we get
$\dfrac{A^{2}}{(k+1+\beta)^{2\delta r}}\geq\dfrac{A^{2}}{(k+\beta)^{2\delta r}}-\dfrac{2\delta rA^{2}}{(k+\beta)^{2\delta r+1}}.$
Moreover, $\tfrac{2\delta rA^{2}}{(k+\beta)^{2\delta r+1}}\leq\tfrac{\tfrac{1}{2}\lambda\tau A^{\tfrac{1}{\delta}}}{(k+\beta)^{2r}}$ if and only if $\big{(}\tfrac{4\delta rA^{2-\tfrac{1}{\delta}}}{\lambda\tau}\big{)}^{\tfrac{1}{1-2r(1-\delta)}}\leq k+\beta.$ Thus, 4.14 holds as
$\big{(}\tfrac{4r\delta A^{\tfrac{2\delta-1}{\delta}}}{\lambda\tau}\big{)}^{\tfrac{1}{1-2r(1-\delta)}}\leq\beta,$
giving our desired results. ∎
Theorem 4.8 implies that the subgradient methods projected with decreasing step size $\alpha_{k}=\lambda(k+\beta)^{-r}$ when $10<r<$ 1 have a convergence rate of $O(K^{-r\delta})$. However, for the case where $\tfrac{1}{\nu+1}<\delta<1$ we can achieve the convergence rate of $O\Big{(}K^{\tfrac{-\delta}{2(1-\delta)}}\Big{)}$. Our next result establishes the sublinear convergence of the projected subgradient method with diminishing step-size $\alpha_{k}=\lambda(k+\beta)^{-r}$ with $r>1$.
Theorem 4.9 (Convergence rate of diminishing method).
Let $\tfrac{1}{\nu+1}<\delta<1$, $r:=\tfrac{1}{2(1-\delta)}$, $A:=(\tfrac{8\delta r}{\tau^{2}})^{\delta r}$, and $\lambda:={\tfrac{1}{2}}\tau A^{\tfrac{1}{\delta}}$ with $\tau=\tfrac{\mu}{L}$, let the sequence $\{x_{k}\}_{k\geq 0}$ be generated by Algorithm 1 with step-size $\alpha_{k}=\lambda(k+\beta)^{-r}$, and let the initial point $x_{0}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$ satisfying $\operatorname{dist}(x_{0};\mathcal{X}^{*})\leq A\beta^{-\delta r},$ with $\beta=\max{\mathopen{}\left\{4\delta r,A^{\tfrac{1}{\delta r}}(\tfrac{2\rho}{\mu})^{\tfrac{2(1-\delta)}{\delta(1+\nu)-1}}\right\}\mathclose{}}$. Then,
$\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq\dfrac{A}{(k+\beta)^{\delta r}}.$ (4.15)
Proof.
By induction, we show the inequality 4.15. The result for $k=0$ holds. We now assume that the inequality 4.15 holds for iteration $k$. We will show that it remains true for $k+1$. First, we observe that
$\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq\frac{A}{(k+\beta)^{\delta r}}\leq\frac{A}{\beta^{\delta r}}\leq\frac{A}{A(\tfrac{2\rho}{\mu})^{\tfrac{2(1-\delta)\delta r}{\delta(1+\nu)-1}}}=(\tfrac{\mu}{2\rho})^{\tfrac{\delta}{\delta(1+\nu)-1}},$
i.e., $x_{k}\in\mathcal{T}_{\gamma}$. Thanks to Lemma 4.1, it holds that
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\lambda(k+\beta)^{-r}\tau\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})+\lambda^{2}(k+\beta)^{-2r}.$ (4.16)
We now consider the convex function $\varphi(t):=t-\lambda(k+\beta)^{-r}\tau t^{\tfrac{1}{2\delta}}+\lambda^{2}(k+\beta)^{-2r}$ on $[0,A^{2}(k+\beta)^{-2\delta r}]$, leading to
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\varphi(\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*}))\leq\max{\mathopen{}\left\{\varphi(0),\varphi(A^{2}(k+\beta)^{-2\delta r})\right\}\mathclose{}}$
$=\max{\mathopen{}\left\{\lambda^{2}(k+\beta)^{-2r},\leavevmode\nobreak\ A^{2}(k+\beta)^{-2\delta r}-\lambda\tau A^{\tfrac{1}{\delta}}(k+\beta)^{-2r}+\lambda^{2}(k+\beta)^{-2r}\right\}\mathclose{}},$
where 4.16 was applied in the first inequality. To establish the desired result, it suffices to show
$\hskip 28.45274pt\dfrac{\lambda^{2}}{(k+\beta)^{2r}}\leq\dfrac{A^{2}}{(k+1+\beta)^{2\delta r}},$ (4.17)
$\dfrac{A^{2}}{(k+\beta)^{2\delta r}}+\dfrac{\lambda^{2}-\lambda\tau A^{\tfrac{1}{\delta}}}{(k+\beta)^{2r}}\leq\dfrac{A^{2}}{(k+1+\beta)^{2\delta r}}.$ (4.18)
Let us begin with verifying 4.17. Inasmuch as the function $\frac{x^{2r}}{(1+x)^{2\delta r}}$ is increasing on $[0,\infty)$,
$\dfrac{\beta^{2r}}{(1+\beta)^{2\delta r}}\leq\dfrac{(k+\beta)^{2r}}{(k+1+\beta)^{2\delta r}}.$ (4.19)
Using the convexity of the function $\frac{1}{x^{2\delta r}}$ on the positive real numbers, we get
$\dfrac{1}{(1+\beta)^{2\delta r}}\geq\dfrac{1}{\beta^{2\delta r}}-\dfrac{2\delta r}{\beta^{2\delta r+1}}=\dfrac{1}{\beta^{2\delta r}}-\dfrac{2\delta r}{\beta^{2r}},$ (4.20)
with $2\delta r+1=2r$. Combining 4.19 with 4.20, it follows that
$\frac{(k+\beta)^{2r}}{(k+1+\beta)^{2\delta r}}\geq\frac{\beta^{2r}}{(1+\beta)^{2\delta r}}\geq\frac{\beta^{2r}}{\beta^{2\delta r}}-\frac{2\delta r\beta^{2r}}{\beta^{2r}}=\beta-2\delta r\geq 4\delta r-2\delta r=2\delta r=\frac{\lambda^{2}}{A^{2}},$
ensuring 4.17. We next justify the inequality 4.18. From the convexity of $\dfrac{A^{2}}{x^{2\delta r}}$, we obtain
$\frac{A^{2}}{(k+1+\beta)^{2\delta r}}-\frac{A^{2}}{(k+\beta)^{2\delta r}}\geq\frac{-2\delta rA^{2}}{(k+\beta)^{2\delta r+1}}=\frac{-2\delta rA^{2}}{(k+\beta)^{2r}}.$
By combining this inequality with $\lambda^{2}-\lambda\tau A^{\tfrac{1}{\delta}}=-2\delta rA^{2}$, which follows from the definition of $A$, $\lambda$, and $r$, the inequality 4.18 is verified. ∎
We next focus on a geometrically decaying step-size to ensure linear convergence to the optimal solution set. In this case, the step-size forms a decreasing geometric sequence that tends to zero, i.e., $\alpha_{k}=\lambda q^{k}$ where $\lambda>0$ and $10<q<$ 1. Theorem 4.10 establishes the linear convergence of the projected subgradient method with geometrically decaying step-size for the case $\delta=1$.
Theorem 4.10 (Convergence rate of geometrically decaying method).
Let $\delta=1$, $\tau=\tfrac{\mu}{L}\neq 1$ and
$\max\Big{\{}\tfrac{5\tau^{2}-4}{2\tau^{2}},0\Big{\}}<\gamma<{\tfrac{1}{2}},\quad 0<\lambda<\tfrac{\tau}{2}{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{1}{\nu}},\quad q:=\sqrt{1-\tfrac{(1-2\gamma)\tau^{2}}{4}},\quad A:=\max{\mathopen{}\left\{\tfrac{2\lambda}{\tau},\operatorname{dist}(x_{0};\mathcal{X}^{*})\right\}\mathclose{}}.$
Let the sequence $\{x_{k}\}_{k\geq 0}$ be generated by Algorithm 1 with step-size $\alpha_{k}=\lambda q^{k}$ and let the initial point $x_{0}\in\mathcal{T}_{\gamma}$ with $\gamma=\nicefrac{{1}}{{2}}$ satisfying
$\operatorname{dist}(x_{0};\mathcal{X}^{*})\leq\tfrac{2\lambda}{\tau-\sqrt{\tau^{2}-4(1-q^{2})}}.$ (4.21)
Then,
$\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq Aq^{k}.$ (4.22)
Proof.
By Lemma 3.19, it follows that $\tau\in(0,1)$. Consequently, $\tfrac{5\tau^{2}-4}{2\tau^{2}}<{\tfrac{1}{2}},$ which ensures the existence of $\gamma$ within the specified range. Additionally, it is evident that $10<q<$ 1. Observing that $1-q^{2}<\tfrac{\tau^{2}}{4}$, we deduce $\tau^{2}-4(1-q^{2})>0$, i.e., the inequality 4.21 is well-defined. We proceed by induction to verify the inequality 4.22. For $k=0$, we note that $\operatorname{dist}(x_{0};\mathcal{X}^{*})\leq A=Aq^{0}$. Assume that 4.22 holds for $i=0,1,\ldots,k$, and we show that it holds for $k+1$. Using the upper bound of $\lambda$ and the assumption $x_{0}\in\mathcal{T}_{\gamma}$, we get
$\operatorname{dist}(x_{k};\mathcal{X}^{*})\leq Aq^{k}\leq A=\max{\mathopen{}\left\{\tfrac{2\lambda}{\tau},\operatorname{dist}(x_{0};\mathcal{X}^{*})\right\}\mathclose{}}\leq{\mathopen{}\left(\tfrac{\mu}{2\rho}\right)\mathclose{}}^{\tfrac{1}{\nu}}.$
Hence, $x_{k}\in\mathcal{T}_{\gamma}$, i.e., applying Lemma 4.1 for the case $\delta=1$ yields
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\lambda q^{k}\tau\operatorname{dist}(x_{k};\mathcal{X}^{*})+\lambda^{2}q^{2k}.$
Let us define the convex function $\varphi(t):=t^{2}-\lambda q^{k}\tau t+\lambda^{2}q^{2k}$ on $[0,Aq^{k}]$, which attains its maximum at 00 or $Aq^{k}$. By the inductive assumption, $\operatorname{dist}(x_{k};\mathcal{X}^{*})\in[0,Aq^{k}]$, i.e.,
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\varphi\big{(}\operatorname{dist}(x_{k};\mathcal{X}^{*})\big{)}\leq\max{\mathopen{}\left\{\varphi(0),\varphi(Aq^{k})\right\}\mathclose{}}=q^{2k}\max{\mathopen{}\left\{\lambda^{2},\leavevmode\nobreak\ A^{2}-\lambda\tau A+\lambda^{2}\right\}\mathclose{}}.$
To ensure $\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq A^{2}q^{2(k+1)}$, it is sufficient to verify
$\lambda^{2}\leq A^{2}q^{2},\quad A^{2}-\lambda\tau A+\lambda^{2}\leq A^{2}q^{2}.$ (4.23)
For the first inequality, since $A^{2}\geq\tfrac{4\lambda^{2}}{\tau^{2}}$ and $\tau^{2}\leq q^{2}$, it follows that $\lambda^{2}\leq A^{2}q^{2}$, where $\tau^{2}\leq q^{2}$ arises from the lower bound of $\gamma$. To show the second inequality in 4.23, we consider the quadratic equation
${\mathopen{}\left(1-q^{2}\right)\mathclose{}}Z^{2}-\lambda\tau Z+\lambda^{2}=0,$ (4.24)
where its positive roots, $z_{1}\leq z_{2}$ are given by
$z_{1,2}:=\tfrac{\lambda\tau\pm\lambda\sqrt{\tau^{2}-4(1-q^{2})}}{2(1-q^{2})}=\tfrac{2\lambda}{\tau\mp\sqrt{\tau^{2}-4(1-q^{2})}}.$
To satisfy the second inequality in 4.23, it sufficient to verify that $z_{1}\leq A\leq z_{2}$. Regarding the smaller root, it is clear that $z_{1}\leq\tfrac{2\lambda}{\tau}\leq A$. For the larger root, we consider two cases: (i) $A=\operatorname{dist}(x_{0};\mathcal{X}^{*})$; (ii) $A=\tfrac{2\lambda}{\tau}$. In Case (i), $A=\operatorname{dist}(x_{0};\mathcal{X}^{*})$, the upper bound in 4.21 and $10<q<$ 1 ensure the result. In Case (ii), $A=\tfrac{2\lambda}{\tau}$, it is clear that $A\leq\tfrac{2\lambda}{\tau-\sqrt{\tau^{2}-4(1-q^{2})}}=z_{2}.$ Therefore, the inductive step is complete, adjusting the result. ∎
4.3 Projected subgradient method with Scaled Polyak’s step-size
In this section, we concentrate on the projected subgradient method with Scaled Polyak’s step-size, $\alpha_{k}=\tfrac{f(x_{k})-f^{*}}{\sigma\|\zeta_{k}\|}$ where $\sigma>0$. Unlike the projected subgradient methods discussed in the previous sections, this approach requires knowing the optimal value of Problem 1.1. This requirement is reasonable in certain scenarios, such as when employing the exact penalty method to solve nonlinear equations. We conclude this section by analyzing the convergence of the sequence generated by Algorithm 1 with a scaled version of Polyak’s step-size. We will show that the sequence generated by this method attains the sublinear convergence if $\tfrac{1}{1+\nu}<\delta<1$, and enjoys the Q-linear convergence if $\delta=1$.
Theorem 4.11 (Convergence rate of scaled Polyak’s method).
Fix the parameters $\sigma>\tfrac{1}{2}$ and $0<\gamma<{\tfrac{2\sigma-1}{2\sigma}}$. Let the iterations $\{x_{k}\}_{k\geq 0}$ be generated by Algorithm 1 with Scaled Polyak’s step-size and initial point $x_{0}\in\mathcal{T}_{\gamma}$. Then, the following assertions hold:
  • (a)
    For each integer $k\geq 0$, $x_{k}\in\mathcal{T}_{\gamma}$ and
    $\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\bigg{(}1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau_{k}^{2}\bigg{)}\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*}),$ (4.25)
    where $\tau_{k}:=\tau\operatorname{dist}^{\tfrac{1}{\delta}-1}(x_{k};\mathcal{X}^{*})$ and $\tau=\tfrac{\mu}{L}$.
  • (b)
    The sequence $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ either converges to zero in a finite number of steps or converges to zero monotonically and
    $\lim_{k\to\infty}f(x_{k})=f^{*}.$
  • (c)
    If $\delta=1$, the sequence $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ Q-linearly converges with rate $\sqrt{1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau^{2}}$ where $\tau\in(0,1)$.
  • (d)
    If $\tfrac{1}{1+\nu}<\delta<1$, the sequence $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ sublinearly converges, i.e., there exists $\eta>0$ such that
    $\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})\leq\eta k^{-\tfrac{\delta}{1-\delta}},$
    for all $k\in\mathbb{N}$ sufficiently large.
  • (e)
    For each $k\in\mathbb{N}$,
    $f^{*}_{k}-f^{*}\leq\frac{\sigma L\operatorname{dist}(x_{0};\mathcal{X}^{*})}{\sqrt{2\sigma(1-\gamma)-1}\sqrt{k+1}},$ (4.26)
    where $f^{*}_{k}:=\min\{f(x_{1}),f(x_{2}),\ldots,f(x_{k})\}$.
Proof.
We first show that $x_{k}\in\mathcal{T}_{\gamma}$ using induction. We simultaneously establish that the inequality 4.25 holds. For the case $k=0$, we have $x_{k}\in\mathcal{T}_{\gamma}$ by assumption. Assume that $x_{k}\in\mathcal{T}_{\gamma}$ for some $k\geq 0$ (induction assumption). We prove that $x_{k+1}\in\mathcal{T}_{\gamma}$ for $k=k+1$. To show the inductive step, two cases are recognized: (i) $\zeta_{k}=0$; (ii) $\zeta_{k}\neq 0$. In Case (i), $\zeta_{k}=0$, it follows from the definition of the projected subgradient method and induction assumption that $x_{k+1}=x_{k}\in\mathcal{T}_{\gamma}$. Moreover, $0=\zeta_{k}\in\partial f(x_{k})$ and $x_{k}$ is a stationary point for problem 1.1. Hence, $x_{k}\in\mathcal{X}^{*}$ by Lemma 3.18. Thus, $\operatorname{dist}(x_{k+1};\mathcal{X}^{*})=\operatorname{dist}(x_{k};\mathcal{X}^{*})=0$, and 4.25 is satisfied. In Case (ii), $\zeta_{k}\neq 0$, let us further consider two subcases: either $x_{k}\in\mathcal{X}^{*}$ or $x_{k}\not\in\mathcal{X}^{*}$. In the case $x_{k}\in\mathcal{X}^{*}$, it follows that $f(x_{k})=f^{*}$, $\alpha_{k}=0$, $\operatorname{dist}(x_{k};\mathcal{X}^{*})=0$, and $\tau_{k}=0$. Then, we obtain $x_{k+1}=\operatorname{proj}_{X}{(x_{k})}=x_{k}\in\mathcal{T}_{\gamma}$, and also 4.25 holds.
Now, let us examine the latter case $x_{k}\not\in\mathcal{X}^{*}$. We choose $\bar{x}\in\operatorname{proj}_{\mathcal{X}^{*}}{(x_{k})}$ implying $\|x_{k}-\bar{x}\|=\operatorname{dist}(x_{k};\mathcal{X}^{*})$ and $f(\bar{x})=f^{*}$. Drawing on the nonexpansiveness of the projection operator and Proposition 3.13, we have
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\|x_{k+1}-\bar{x}\|^{2}\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{2\big{(}f(x_{k})-f(\bar{x})\big{)}}{\sigma\|\zeta_{k}\|^{2}}\langle\zeta_{k},\bar{x}-x_{k}\rangle+\tfrac{\big{(}f(x_{k})-f(\bar{x})\big{)}^{2}}{\sigma^{2}\|\zeta_{k}\|^{2}}\vspace{2mm}$
$\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{2\big{(}f(x_{k})-f(\bar{x})\big{)}}{\sigma\|\zeta_{k}\|^{2}}\Big{(}f(\bar{x})-f(x_{k})+\rho\|x_{k}-\bar{x}\|^{\nu+1}\Big{)}+\tfrac{\big{(}f(x_{k})-f(\bar{x})\big{)}^{2}}{\sigma^{2}\|\zeta_{k}\|^{2}}\vspace{2mm}$
$=\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{f(x_{k})-f^{*}}{\sigma^{2}\|\zeta_{k}\|^{2}}\Big{(}2\sigma\rho\operatorname{dist}^{\nu+1}(x_{k};\mathcal{X}^{*})-(2\sigma-1)(f(x_{k})-f^{*})\Big{)}.$ (4.27)
Since $x_{k}\in\mathcal{T}_{\gamma}\subseteq\mathcal{T}_{1}$, we get $\operatorname{dist}(x_{k};\mathcal{X}^{*})<{\mathopen{}\left(\gamma\tfrac{\mu}{\rho}\right)\mathclose{}}^{\tfrac{\delta}{\delta(1+\nu)-1}}$ and $\|\zeta_{k}\|\leq L$. Therefore,
$\rho\operatorname{dist}^{\nu+1}(x_{k};\mathcal{X}^{*})<\gamma\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})\leq\gamma(f(x_{k})-f^{*}),$ (4.28)
where the second inequality comes from Hölderian property of $f$. Applying these inequalities in 4.27, we deduce
$\begin{array}[]{ll}\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})&\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{f(x_{k})-f^{*}}{\sigma^{2}\|\zeta_{k}\|^{2}}\Big{(}2\sigma\gamma\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})-(2\sigma-1)\mu\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})\Big{)}\vspace{2mm}\\ &\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{\mu^{2}\operatorname{dist}^{\tfrac{2}{\delta}}(x_{k};\mathcal{X}^{*})}{\sigma^{2}\|\zeta_{k}\|^{2}}\Big{(}2\sigma\gamma-2\sigma+1\Big{)}\leq\bigg{(}1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau_{k}^{2}\bigg{)}\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*}).\end{array}$
Hence, the inequality 4.25 is valid. Additionally, inasmuch as $x_{k}\in\mathcal{T}_{\gamma}$ and $x_{k}\not\in\mathcal{X}^{*}$, we conclude $\tau_{k}\in(0,1]$ by Lemma 3.19, i.e., $10<1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau_{k}^{2}<$ 1. Hence, $\operatorname{dist}(x_{k+1};\mathcal{X}^{*})<\operatorname{dist}(x_{k};\mathcal{X}^{*})$, which yields $x_{k+1}\in\mathcal{T}_{\gamma}$.
In order to prove the convergence of $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ (Assertion (ii)), based on Assertion (i) either there exists some $k_{0}\in\mathbb{N}$ such that $\operatorname{dist}(x_{k};\mathcal{X}^{*})=0$ for all $k\geq k_{0}$, or $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ is a decreasing positive sequence. We show that $\operatorname{dist}(x_{k};\mathcal{X}^{*})\to 0$ in the latter case. Suppose for contradiction that $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ converges to some $d>0$. Taking 4.25 into account, we get $d^{2}\leq\bigg{(}1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau^{2}d^{\tfrac{2}{\delta}-2}\bigg{)}\leavevmode\nobreak\ d^{2}$, leading to a contradiction. Hence, $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ decreasingly converges to zero. Moreover, the closedness of $\mathcal{X}^{*}$ implies that $x_{k}\to x^{*}$, for some $x^{*}\in\mathcal{X}^{*}$, i.e., it results in $f(x_{k})\to f(x^{*})=f^{*}$, due to continuity of $f$.
Let us verify Assertion (iii) for $\delta=1$. In this case, Lemma 3.19 ensures that $\tau_{k}=\tau=\tfrac{\mu}{L}\in(0,1]$. It follows from inequality 4.25 that
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\leq\bigg{(}1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau^{2}\bigg{)}\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*}),$
where $10<1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau^{2}<$ 1. Therefore, $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ converges Q-linearly to zero with rate $\sqrt{1-\Big{(}2\sigma(1-\gamma)-1\Big{)}\sigma^{-2}\tau^{2}}$. Consequently, $x_{k}$ converges $\mathcal{X}^{*}$.
We investigate Assertion (iv) for $\tfrac{1}{1+\nu}<\delta<1$. The inequality 4.25 can be rewritten as
$\operatorname{dist}^{\tfrac{1}{\delta}}(x_{k};\mathcal{X}^{*})\leq\dfrac{\sigma^{2}}{\Big{(}2\sigma(1-\gamma)-1\Big{)}\tau^{2}}\Big{(}\operatorname{dist}(x_{k};\mathcal{X}^{*})-\operatorname{dist}(x_{k+1};\mathcal{X}^{*})\Big{)}.$
Thus, the sublinear convergence of $\operatorname{dist}(x_{k};\mathcal{X}^{*})$ is guaranteed from Fact 2.1.
To show Assertion (v), combining 4.27 and 4.28 with $\|\zeta_{k}\|\leq L$, it follows that
$\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})$ $\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})+\tfrac{f(x_{k})-f^{*}}{\sigma^{2}\|\zeta_{k}\|^{2}}\Big{(}2\sigma\gamma(f(x_{k})-f^{*})-(2\sigma-1)(f(x_{k})-f^{*})\Big{)}$
$\leq\operatorname{dist}^{2}(x_{k};\mathcal{X}^{*})-\tfrac{(f(x_{k})-f^{*})^{2}}{\sigma^{2}L^{2}}\Big{(}2\sigma(1-\gamma)-1\Big{)}\leq\ldots\leq\operatorname{dist}^{2}(x_{0};\mathcal{X}^{*})-\tfrac{2\sigma(1-\gamma)-1}{\sigma^{2}L^{2}}\sum_{i=0}^{k}(f(x_{i})-f^{*})^{2}$
$\leq\operatorname{dist}^{2}(x_{0};\mathcal{X}^{*})-\tfrac{2\sigma(1-\gamma)-1}{\sigma^{2}L^{2}}(f^{*}_{k}-f^{*})^{2}(k+1).$
Since $\operatorname{dist}^{2}(x_{k+1};\mathcal{X}^{*})\geq 0$, the inequality 4.26 is obtained, adjusting our claims. ∎
5 Application to robust low-rank matrix recovery
This section presents a detailed analysis of our numerical experiments to demonstrate the efficiency of various projected subgradient methods employing several step-size strategies, including the following:
  • $\bullet$
    Diminishing: Algorithm 1 with the step-size $\alpha_{k}:=\frac{\alpha_{0}}{\sqrt{k+1}}$ and $\alpha_{0}>0$;
  • $\bullet$
    Decaying: Algorithm 1 with the step-size $\alpha_{k}:=\alpha_{0}\lambda^{k}$ and $\alpha_{0}>0$;
  • $\bullet$
    Polyak: Algorithm 1 with the step-size $\alpha_{k}:=\tfrac{f(x_{k})-f^{*}}{\|\zeta_{k}\|}$;
  • $\bullet$
    Scaled Polyak: Algorithm 1 with the step-size $\alpha_{k}:=\tfrac{f(x_{k})-f^{*}}{4\|\zeta_{k}\|}$.
The experiments are performed on various applications of the robust low-rank matrix recoveries using different datasets. The applications include robust matrix completion on the MovieLens dataset, image inpainting, robust nonnegative matrix factorization for face recognition, and matrix compression. The base of all these applications is the robust matrix recovery problem
$\min_{U,V}\|X-UV\|_{1},$ (5.1)
where the matrix $X$ is approximated by $UV$. Here, $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{r\times n}$ represent low-rank matrices li2020nonconvex . This formulation results in a nonconvex and nonsmooth paraconvex optimization problem (as confirmed in Proposition 3.14), which accepts significant computational challenges. In our implementation, the code was entirely written in Python (a $3.11.7$ version) and tested on a MacBook Pro, macOS Ventura $13.6.6$, equipped with 32 GB of RAM.
5.1 Robust matrix completion problem
In data mining and machine learning, many datasets are represented as matrices, often with missing values ranging from minimal to substantial. The problem of robust matrix completion (RMC) wang2012probabilistic is formulated as
$\min_{U\in\mathbb{R}^{m\times r},V\in\mathbb{R}^{r\times n}}\frac{1}{2}\|M\odot(X-UV)\|_{1},$ (5.2)
where $\odot$ stands for the Hadamard product of two matrices. Missing values can be addressed by defining a binary weight matrix $M$ given by
$M_{ij}={\mathopen{}\left\{\begin{array}[]{ll}1&\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{if}\leavevmode\nobreak\ X_{ij}\leavevmode\nobreak\ \mathrm{is\leavevmode\nobreak\ known},\\ 0&\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{if}\leavevmode\nobreak\ X_{ij}\leavevmode\nobreak\ \mathrm{is\leavevmode\nobreak\ unknown}.\end{array}\right.\mathclose{}}$
Recovering such matrices is challenging due to sparsity in observed values, compounded by noise, corruption, and outliers.
We evaluated our method on the widely-used MovieLens (ML-100K) dataset, which contains approximately 100,000 ratings from 943 users for 1,682 movies (ratings range from 1 to 5). The dataset was split into an $80\%$ training set and a $20\%$ test set to evaluate reconstruction quality. The proposed algorithms are applied to the problem (5.2) with the rank $r=20$ and a maximum of $1000$ iterations. For the Decaying strategy, the parameters are set as $\gamma_{k}=0.95$ and $\alpha_{0}=10^{-3}$, while for the Diminishing strategy, $\alpha_{0}=10^{-3}$. The reconstruction accuracy is measured by the root mean squared error (RMSE), which the results are presented in Tabel 1, and the convergence behavior is analyzed based on the loss values iterations, where the results are presented in Figure 3. To demonstrate the robustness of our algorithm, we used two different initialization methods: one based on SVD (shown in the left subfigure of Figure 3) and the other with a random starting point (shown in the right subfigure of Figure 3).
Figure 3 illustrates that the Polyak and Scaled Polyak strategies achieved lower loss values compared to the Decaying and Diminishing strategies under both initialization methods. However, Table 3 indicates the Decaying strategy achieved the best performance with an RMSE of $1.153$, closely followed by Scaled Polyak with an RMSE of $1.254$. Both methods indicate fast convergence and accurate predictions. In contrast, the Diminishing algorithm results in the highest RMSE of $1.656$.
Step-size Polyak Scaled Polyak Diminishing Decaying
svd initialization 1.460 1.254 1.655983 1.153
random initialization 1.350 1.288 1.085 1.066
Table 1: RMSE values for the MovieLens 100k dataset.
Figure 3: Convergence of the loss function (log scale) over 1000 iterations for the MovieLens 100K dataset, comparing four step-size strategies: Polyak, Scaled Polyak, Diminishing, and Decaying using two initial points.
5.2 Image inpainting problem
Image inpainting aims at reconstructing missing or corrupted regions of an image, aiming to restore it as closely as possible to its original form; cf. guillemot2013image . Here, we applied robust matrix completion (RMC) to images corrupted with $40\%$ noise, where $40\%$ of the pixel values were masked. The experimental setup uses the following parameters: $r=100$, $iterations=1000$, and $\alpha_{0}=10^{-2}$ for diminishing step-size and $\alpha_{0}=10^{-3}$ for decaying step-size. The quality of reconstruction is assessed using the peak signal-to-noise ratio (PSNR) and the signal-to-noise ratio (SNR), summarized in Table 2, and loss convergence in Figure 6. The main images with their $40\%$ corrupted are plotted in Figure 4, and the reconstructed ones are plotted in Figure 5.
Image Polyak Scaled Polyak Diminishing Decaying
PSNR SNR PSNR SNR PSNR SNR PSNR SNR
lighthouse 20.86 7.39 21.25 7.78 17.64 4.17 15.06 1.60
man 25.99 13.12 26.47 13.60 24.22 11.34 20.91 8.04
kiel 23.04 8.11 21.32 6.40 18.05 3.13 15.10 0.18
houses 21.76 9.77 20.96 8.97 16.50 4.51 14.05 2.06
Table 2: Quantitative evaluation of image reconstruction using the PSNR and SNR for four images (Lighthouse, Man, Kiel, and Houses), corrupted with $40\%$ noise. The results are reported for four step-size strategies: Polyak, Scaled Polyak, Diminishing, and Decaying. All values are rounded to two decimals.
Figure 4: The first row displays the original images: "Man" (a), "Kiel" (b), "Lighthouse" (c), and "Houses" (d). The second row shows the corresponding corrupted images (e–h) with $40\%$ random uniform noise applied.
The results in Figure 5, Figure 6, and Table 2 show that the Polyak and Scaled Polyak strategies consistently outperformed the Diminishing and Decaying methods across all metrics. Notably, the Scaled Polyak strategy achieved the highest PSNR and SNR values for most images, including the PSNR of 26.47 dB and the SNR of 13.60 for the ’Man’ image, indicating superior reconstruction accuracy. The Polyak strategy also delivered strong results, with the PSNR of 25.99 dB and the SNR of 13.12 for the same image. In contrast, the Diminishing and Decaying methods demonstrated significantly poorer performance. The convergence plots in Figure 6 highlight that Scaled Polyak and Polyak strategies provide the fastest and smoothest loss reductions, outperforming the Diminishing and Decaying methods. The Decaying strategy showed minimal progress after an initial decrease, while the Diminishing strategy exhibited slow and inconsistent convergence. The Scaled Polyak strategy achieved the best convergence rates, which is closely followed by the Polyak strategy.
Figure 5: Qualitative comparison of image inpainting results for four corrupted images (lighthouse, man, kiel, and houses) using Polyak, Scaled Polyak, Diminishing, and Decaying step-size strategies. Each column shows the reconstructions for the step-size method.
Figure 6: Qualitative comparison of image inpainting results for four corrupted images (Lighthouse, Man, Kiel, and Houses) using Polyak, Scaled Polyak, Diminishing, and Decaying step-size strategies. Each column shows the reconstructions for the step-size method.
5.3 Matrix compression using RNMF
Matrix compression reduces large matrices’ storage and computational demands by approximating them with compact representations that retain essential information while minimizing redundancy; cf. yuan2005projective . This process is critical in fields like data analysis, machine learning, and scientific computing, where efficiently handling high-dimensional data is essential. Techniques such as low-rank approximations, including nonnegative matrix factorization (NMF), are widely used. These methods decompose the original matrix into two lower-rank factors that, when combined, approximate the original data, enabling efficient storage, faster computations, and, in some cases, enhanced interpretability of the underlying structures.
Here, we applied the projected subgradient method to the reformulated robust matrix completion (RMC) problem in (5.1), treating it as a robust nonnegative matrix factorization (RNMF) problem
$\min_{U\geq 0\in\mathbb{R}^{m\times r},V\geq 0\in\mathbb{R}^{r\times n}}\frac{1}{2}\|X-UV\|_{1},$ (5.3)
using the aforementioned step-size strategies on the ’Man’ image. The goal is to compress the image by representing it with two low-rank matrices for different ranks ($r=5,10,50,100$).
The results in Figures 7 and 8 illustrate the impact of rank on reconstruction quality during matrix compression. Notably, the reconstruction quality is improved across all step-size strategies as the rank increased from 5 to 100. For lower ranks (5 and 10), the Decaying and Diminishing methods are struggling, leaving significant visual artifacts in the reconstructions. In contrast, the Scaled Polyak and Polyak methods demonstrate better stability and higher-quality reconstructions, even at lower ranks. All methods improved at higher ranks (50 and 100), but the Scaled Polyak method consistently produced sharper and more accurate results. Figure 8 further highlights the convergence behavior of the step-size strategies across ranks. The Decaying method stagnates early and fails to reduce the loss significantly. The Diminishing method shows some progress but at a slower rate. However, the Polyak and Scaled Polyak methods effectively reduce the loss, with Scaled Polyak consistently achieving the lowest final loss values. These findings show that while rank significantly affects reconstruction performance, the Polyak-based methods remain robust across different ranks, making them particularly effective for matrix compression tasks.
Figure 7: Comparison of reconstructed images for different ranks ($r=5,10,50,100$) using four step-size strategies: Decaying, Diminishing, Polyak, and Scaled Polyak.
Figure 8: Convergence plots showing the loss value (log scale) over 1000 iterations for different ranks ($r=5,10,50,100$) using four step-size strategies: Decaying, Diminishing, Polyak, and Scaled Polyak.
5.4 Face recognition using RNMF
NMF has been extensively applied to face recognition tasks due to its ability to decompose facial images into additive, parts-based representations that align with human facial features’ intuitive understanding; cf. guillamet2002non . NMF captures localized features such as eyes, noses, and mouths by factorizing a nonnegative data matrix into two nonnegative matrices, facilitating more interpretable and efficient data representations.
We here applied the projected subgradient methods to the RNMF (5.3) problem on the Olivetti Faces dataset for dimensionality reduction. The performance is evaluated using a K-Nearest Neighbors (KNN) classifier for three different neighborhood sizes ($k=1,3,5$). For the reconstruction process, we set the rank $r=128$ and a maximum of iterations to $1000$. The reconstructed images are assessed as quantitative metrics. Two initialization methods were used: Nonnegative SVD-R trigeorgis2014deep (Figure 10(a)) and Nonnegative random initialization (Figure 10(b)). The results indicate that the choice of starting point plays a crucial role in RNMF performance, as highlighted in Figure 10. All methods performed well for $k=1$ with Polyak achieving the highest accuracy, $92.5\%$. For ($k=3$), Polyak again shows the best accuracy ($86.2\%$), followed by Scaled Polyak and Diminishing ($85.0\%$), while Decaying dropped to $82.5\%$. For $k=5$, Polyak and Diminishing both achieve ($85.0\%$), Scaled Polyak slightly lags at ($82.5.0\%$) and Decaying dropped further to ($78\%$). These findings suggest that while Polyak and Diminishing show stability across different neighborhood sizes, Scaled Polyak tends to have slightly reduced performance as complexity increases, and Decaying demonstrates clear limitations in preserving discriminative features during dimensionality reduction. Overall, Polyak’s step-size strategy stands out for its superior generalization capability in high-dimensional data scenarios, making it a robust choice for complex classification tasks.
Figure 9: Visual comparison of original and reconstructed images from the Olivetti Faces dataset using four step-size strategies: Polyak, Scaled Polyak, Decaying, and Diminishing. Each subfigure displays original images (left) alongside their reconstructions (right) for 40 selected individuals.
Figure 10: Loss convergence (log scale) for different step-size strategies on the Olivetti Faces dataset over 1000 iterations with two different starting points.
Step-size Polyak Scaled Polyak Diminishing Decaying
Class 1 0.925 0.913 0.913 0.875
Class 3 0.862 0.850 0.850 0.825
Class 5 0.850 0.825 0.850 0.788
Table 3: Classification accuracy of K-Nearest Neighbors (KNN) for the Olivetti Faces dataset across three neighborhood sizes ($k=1,2,3$) after dimensionality reduction using NMF.
6 Concluding remarks
In this paper, we studied the projected subgradient methods for paraconvex functions under the Hölderian error bound condition. In the first step, we investigate the properties and characterizations of paraconvex functions, including their local Lipschitz continuity. Notably, we established that $\nu$-paraconvexity coincides with $\nu$-weak convexity on convex and compact sets. In addition, under the Hölderian error bound condition, we identified regions around the optimal set that includes no saddle points.
Next, we described the generic projected subgradient method with several step-sizes (i.e., constant, nonsummable diminishing, square-summable but not summable, geometrically decaying, and Scaled Polyak’s step sizes). It was shown that the projected subgradient method with a constant step-size achieves linear convergence up to a fixed threshold. Furthermore, we demonstrated subsequential and global convergence for nonsummable and square-summable but nonsummable step-sizes, respectively, while establishing linear convergence for geometrically decaying and Scaled Polyak’s step-sizes.
Finally, we assessed the practical effectiveness of the proposed methods in robust low-rank matrix recovery problems such as robust matrix completion, image inpainting, matrix compression, and robust nonnegative matrix factorization. Our preliminary numerical experiments highlighted the potential of our methods for nonsmooth and nonconvex problems that lack a suitable structure. Notably, the Scaled Polyak’s step-size consistently outperformed the other considered algorithms, highlighting its practical value and potential for broader adoption in nonconvex optimization applications, e.g., solving high-order and non-Euclidean proximal-point subproblems; cf. ahookhosh2024high ; ahookhosh2021bregman ; kabgani2024high .